### v2

# On the asymptotically linear problem for an elliptic equation with an indefinite nonlinearity

Abstract
We study the semilinear elliptic problem
$\Delta u=Q_{\Omega}|u|^{p-2}u\quad\text{in }\mathbb{R}^{N},$
where $Q_{\Omega}=\chi_{\Omega}-\chi_{\mathbb{R}^{N}\setminus\Omega}$ for a bounded smooth domain $\Omega\subset\mathbb{R}^{N}$, $N\geq 3$, and $1<p<2^{*}$. This equation arises in the study of optical waveguides and exhibits indefinite nonlinearity due to the sign-changing weight $Q_{\Omega}$. We prove that, for $p>2$ sufficiently close to $2$, positive solutions are nondegenerate and the problem admits a unique least energy solution. Our approach combines a detailed analysis of an associated eigenvalue problem involving $Q_{\Omega}$ with variational methods and blow-up techniques in the asymptotically linear regime. We also provide a comprehensive study of the spectral properties of the corresponding linear problem, including the existence and qualitative behavior of eigenfunctions, sharp decay estimates, and symmetry results. In particular, we establish analogues of the Faber–Krahn and Hong–Krahn–Szegö inequalities in this non-standard setting.
Mathematics Subject Classification: 35J61 35A02 (primary) 35J20 35B09 35B40 (secondary).
Keywords: Uniqueness and nondegeneracy, self-focusing core, eigenvalues and eigenfunctions, Faber-Krahn inequality, Asymptotic decay, Symmetry and symmetry breaking
1 Introduction
Let $\Omega$ be a bounded smooth open subset of $\mathbb{R}^{N}$ (not necessarily connected), $N\geq 3$, and let
$Q_{\Omega}(x):=\begin{cases}1&\text{if \ }x\in\Omega,\\ -1&\text{if \ }x\in\mathbb{R}^{N}\smallsetminus\Omega.\end{cases}$
In this paper we study the uniqueness of minimizers and nondegeneracy of positive solutions to the following semilinear problem
$\begin{cases}-\Delta u=Q_{\Omega}|u|^{p-2}u,\\ u\in X^{p}=D^{1,2}(\mathbb{R}^{N})\cap L^{p}(\mathbb{R}^{N}),\end{cases}$ (1.1)
where $1<p<2^{*},$ $2^{*}=\frac{2N}{N-2}$ is the critical Sobolev exponent, and $D^{1,2}(\mathbb{R}^{N})=:\{u\in L^{2^{*}}(\mathbb{R}^{N})\;:\;\nabla u\in L^{2}(\mathbb{R}^{N})\}$. We endow $X^{p}$ with the norm
$\|u\|_{X^{p}}:=(\|u\|^{2}+|u|_{p}^{2})^{\frac{1}{2}},\qquad\text{where \ }\|u\|^{2}:=\int_{\mathbb{R}^{N}}|\nabla u|^{2}$
and $|\cdot|_{p}$ denotes the norm in $L^{p}(\mathbb{R}^{N})$.
Problem (1.1) arises in connection with some models of optical waveguides propagating through a stratified dielectric medium. For a more detailed explanation of the physical relevance of the model, we refer to [23, 24, 1, 14, 9] and the references therein.
It is known that uniqueness does not hold in general for positive solutions to (1.1). In particular, in [7] the authors show that there exists $\varepsilon>0$ such that, if $p\in(2^{*}-\varepsilon,2^{*})$, then multiple positive solutions to (1.1) can be constructed using a Lyapunov-Schmidt procedure, whenever $\Omega$ has nontrivial topology (if it is an annulus, for instance), or if $\Omega$ has a special geometry (e.g. if it is a dumbbell type domain), or if $\Omega$ is disconnected (the union of two disjoint balls, for example). The critical case $p=2^{*}$ is studied in [5], where Lusternik-Schnirelmann theory is used to obtain multiplicity of positive solutions when a sufficiently thin tubular neighborhood of a manifold with rich topology is removed from $\Omega$.
In this paper, we show that the picture is completely different when considering powers $p$ far from the critical exponent and sufficiently close to 2. This is the first result regarding the uniqueness of minimizers (i.e. least energy solutions) and the nondegeneracy of positive solutions to (1.1).
Theorem 1.1.
Let $\Omega$ be a bounded smooth open subset of $\mathbb{R}^{N}$ and $N\geq 3$. There exists $p_{0}>2$ such that the problem (1.1) has a unique positive least energy solution for $p\in(2,p_{0})$. Moreover, all positive solutions of (1.1) are non-degenerate for $p\in(2,p_{0})$.
The exact notion of nondegeneracy used in Theorem 1.1 can be found in Section 3.2 and relates to the fact that the linearized problem at a positive solution has only the trivial solution. The definition of least energy solution can be found in Subsection 3.1
Theorem 1.1 holds for general bounded smooth sets $\Omega$. This contrasts with the case when the exponent $p$ is far from 2, where symmetries on $\Omega$ can be used to obtain multiplicity results. To illustrate this point, we show a symmetry breaking result for least energy solutions in the slightly subcritical regime, namely, if $p$ is close enough to the Sobolev critical exponent $2^{*}=\frac{2N}{N-1}$. By using rotations, this yields the existence of infinitely many different least energy solutions. Recall that foliated Schwarz symmetric functions are axially symmetric functions with some monotonicity properties in the polar angle, we refer to [6], for instance, for a precise definition.
Theorem 1.2.
Let $\Omega:=\{x\in\mathbb{R}^{N}\>:\>\frac{1}{2}<|x|<1\}$ and let $p_{k}\subset(2,2^{\ast})$ be a sequence such that $p_{k}\to 2^{*}$. Then, passing to a subsequence, (1.1) with $p=p_{k}$ has a nonradial foliated Schwarz least energy solution.
The idea behind the proof of Theorem 1.1 is that, in the asymptotically linear case (namely, as $p\to 2$), the solutions inherit the properties of the linear problem ($p=2$), where typically one expects a well-behaved situation. Hence, to prove Theorem 1.1 we need two ingredients: a good knowledge of the linear problem and a good control of the solutions of (1.1) as $p$ tends to 2.
To understand the linear problem, let us consider the eigenvalue problem
$\begin{cases}-\Delta u=\Lambda Q_{\Omega}u,\\ u\in H^{1}(\mathbb{R}^{N}),\quad u\neq 0.\end{cases}$ (1.2)
A solution to this problem is a pair $(\Lambda,u)\in\mathbb{R}\times H^{1}(\mathbb{R}^{N})$ that satisfies (1.2) weakly, i.e., $u\neq 0$ and
$\int_{\mathbb{R}^{N}}\nabla u\cdot\nabla\varphi=\Lambda\int_{\mathbb{R}^{N}}Q_{\Omega}u\varphi\qquad\text{for all \ }\varphi\in C^{\infty}_{c}(\mathbb{R}^{N}).$
This implies, in particular, that
$0<\int_{\mathbb{R}^{N}}|\nabla u|^{2}=\Lambda\int_{\mathbb{R}^{N}}Q_{\Omega}u^{2}.$
Therefore, $\Lambda\neq 0$, and $\int_{\mathbb{R}^{N}}Q_{\Omega}u^{2}>0$ if and only if $\Lambda>0$. The existence of solutions with $\Lambda>0$ was previously established in [25] in a more general framework. However, to the best of our knowledge, a detailed qualitative analysis of these solutions has not been carried out before. In this paper, we provide an ad hoc proof of the existence of solutions to (1.2) and develop new results concerning their structure, including regularity, decay, nodal properties, and symmetry.
To state our main existence result for (1.2), consider the set
$\mathcal{S}_{\Omega}:=\Big\{v\in D^{1,2}(\mathbb{R}^{N}):\int_{\mathbb{R}^{N}}Q_{\Omega}v^{2}=1\Big\}.$ (1.3)
The following result shows that the linear problem (1.2) resembles the case of the Dirichlet Laplacian in bounded domains.
Theorem 1.3.
There is a set $\{\phi_{k}\in\mathcal{S}_{\Omega}:k\in\mathbb{N}\}$ of functions that are orthogonal in $D^{1,2}(\mathbb{R}^{N})$ such that
$\|\phi_{k}\|^{2}=\inf_{v\in V_{k}\cap\mathcal{S}_{\Omega}}\|v\|^{2}=:\Lambda_{k},$
where $V_{1}:=D^{1,2}(\mathbb{R}^{N})$ and $V_{k}$ is the orthogonal complement in $D^{1,2}(\mathbb{R}^{N})$ of the space generated by the eigenfunctions $\{\phi_{1},\ldots,\phi_{k-1}\}$ if $k\geq 2$. Furthermore,
  • $(i)$
    the pair $(\Lambda_{k},\phi_{k})$ solves (1.2),
  • $(ii)$
    $\phi_{k}\in W^{2,p}_{loc}(\mathbb{R}^{N})\cap C^{1,\alpha}_{loc}(\mathbb{R}^{N})$ for all $s\in[1,\infty)$, $\alpha\in(0,1)$ and $k\in\mathbb{N}$,
  • $(iii)$
    $|\phi_{1}|>0$ in $\mathbb{R}^{N}$,
  • $(iv)$
    $0<\Lambda_{1}<\Lambda_{2}\leq\cdots\leq\Lambda_{k}\leq\cdots$ and $\Lambda_{k}\to\infty$. Every eigenspace
[…]
    has finite dimension and $\dim(E_{1})=1$.
Notice that Theorem 1.3 does not state that the sequence $(\phi_{k})$ generates $D^{1,2}(\mathbb{R}^{N})$. This is, in fact, false, due to the admissibility condition $\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{k}^{2}=1$ for all $k\in\mathbb{N}$, see Remark 2.3. The main challenge in the proof of Theorem 1.3 is to control the inherent lack of compactness of equations in unbounded domains, where (1.3) and the boundedness of the $\Omega$ play a crucial role (see [8] a for related problem with $\Omega$ unbounded).
Next, we explore how the semilinear problem (1.1) relates to the linear problem (1.2). For this, we follow the ideas in [21] and introduce an auxiliary nonlinear eigenvalue problem, which is slightly easier to handle variationally. Let
$\mathcal{S}_{\Omega,p}:=\left\{v\in X^{p}:\int_{\mathbb{R}^{N}}Q_{\Omega}(x)|v|^{p}=1\right\},$
and consider the minimization problem
$\alpha_{p}:=\inf_{v\in\mathcal{S}_{\Omega,p}}\|v\|^{2}.$ (1.4)
Theorem 1.4.
For $p\in(1,2^{*})\smallsetminus\{2\}$, there exists a positive function $v_{p}\in\mathcal{S}_{\Omega,p}$ achieving the infimum in (1.4). Moreover, $u_{p}:=\alpha_{p}^{\frac{1}{p-2}}v_{p}$ is a positive solution to (1.1). Furthermore,
  1. 1.
    if $\Lambda_{1}=1$, then
    $u_{p}\to\left(e^{-\frac{1}{2}\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{2}\ln(\phi_{1}^{2})}\right)\phi_{1}\in D^{1,2}(\mathbb{R}^{N})\quad\text{as}\;\;p\to 2,$ (1.5)
    where $\phi_{1}$ is the positive first eigenfunction of (1.2) with $\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{2}=1$,
  2. 2.
    if $\Lambda_{1}>1,$ then
    $\lim_{p\nearrow 2}|u_{p}|_{\infty}=0\qquad\text{and}\qquad\lim_{p\searrow 2}u_{p}(x)=\infty\quad\text{ for a.e. }x\in\mathbb{R}^{N},$
    where $|\cdot|_{\infty}$ is the norm in $L^{\infty}(\mathbb{R}^{N})$, and
  3. 3.
    if $\Lambda_{1}<1,$ then
    $\lim_{p\nearrow 2}u_{p}(x)=\infty\quad\text{ for a.e. }x\in\mathbb{R}^{N}\qquad\text{and}\qquad\lim_{p\searrow 2}|u_{p}|_{\infty}=0.$
Theorems 1.3 and 1.4 are the two main ingredients to prove Theorem 1.1 together with suitable blow-up arguments inspired by [11, 12, 20, 13], where blow-up techniques have been used to show uniqueness (and nondegeneracy) of solutions in the asymptotically linear regime for the Laplacian and the fractional Laplacian in bounded domains. Nevertheless, important obstacles appear due to the sign-changing weight $Q_{\Omega}$, and the classical arguments need to be substantially adapted. In particular, due to these complications, we were not able to show the uniqueness of all positive solutions, but instead we only show the uniqueness of minimizers and leave the uniqueness of positive solutions as an open problem.
We also establish the following sharp decay estimates for solutions to (1.1), whenever $p\in(2,p_{s}]$, where $p_{s}:=\frac{2N-2}{N-2}$ is sometimes called the Serrin exponent.
Proposition 1.5.
Let $p\in(2,p_{s})$ and let $w\in X^{p}$ be a positive solution to $\Delta w=Q_{\Omega}w^{p-1}$ in $\mathbb{R}^{N}$. There is $C>1$ such that,
  1. 1.
    if $p\in(2,p_{s})$,
    $C^{-1}|x|^{-\frac{2}{p-2}}\leq w(x)\leq C|x|^{-\frac{2}{p-2}}\qquad\text{ for all }|x|>1;$ (1.6)
  2. 2.
    if $p=p_{s}=\frac{2N-2}{N-2}$,
    $C^{-1}\left(|x|\sqrt{\ln(|x|)}\right)^{2-N}\leq w(x)\leq C\left(|x|\sqrt{\ln(|x|)}\right)^{2-N}\qquad\text{ for all }|x|>1.$ (1.7)
These sharp estimates answer an open question in [6, Problem 1.6].
We complement our results with some further qualitative properties of the eigenfunctions. We obtain information on their nodal regions, sharp decay, symmetries, Faber-Krahn and Hong–Krahn–Szegö type inequalities, and other results. To be more precise, in the following we use $\phi_{k}$ and $\Lambda_{k}$ to denote the $k$-th eigenfunction and eigenvalue given by Theorem 1.3. In the next result we write explicitly the domain dependency of the eigenvalues, namely, $\Lambda_{k}=\Lambda_{k}(\Omega)$ is the $k$-th eigenvalue associated to (1.2). We also use $\Omega^{*}$ to denote the Schwarz symmetrization of a bounded open set $\Omega$, namely, the unique ball centered at the origin with the same measure as $\Omega$.
Theorem 1.6.
Let $\Omega\subset\mathbb{R}^{N}$ be an open bounded set.
  1. 1.
    (Courant’s result) For every $k\geq 2$, $\phi_{k}$ changes sign and has at most $k$ nodal domains.
  2. 2.
    (Sharp decay) For each $k\in\mathbb{N}$ there is $C>0$ such that
    $|\phi_{k}|\leq C|x|^{-\frac{N-1}{2}}e^{-\sqrt{\Lambda_{k}}|x|}\qquad\text{for all \ }|x|\geq 1.$
    This estimate is sharp, in the sense that there is $c>0$ such that
    $|\phi_{1}|>c|x|^{-\frac{N-1}{2}}e^{-\sqrt{\Lambda_{1}}|x|}\qquad\text{for all \ }|x|\geq 1.$
  3. 3.
    (Faber-Krahn inequality)
    $\Lambda_{1}(\Omega)\geq\Lambda_{1}(\Omega^{*}).$
  4. 4.
    (Hong–Krahn–Szegö inequality)
    $\Lambda_{2}(\Omega)>\Lambda_{1}(2^{-N}\Omega^{*}).$ (1.8)
    This bound is sharp, in the sense that, for every $c>0$,
    $\inf\{\Lambda_{2}(U)\,:\,U\subset\mathbb{R}^{N}\;\text{open bounded such that }|U|=c\,\}=\Lambda_{1}(B_{r}(0)),\qquad r:=\left(\frac{c}{2|B_{1}(0)|}\right)^{\frac{1}{N}},$
    and a minimizing sequence is given by $\Omega_{n}:=B_{r}(-ne_{1})\cup B_{r}(ne_{1})$ for $n\in\mathbb{N},$ where $e_{1}:=(1,0,\ldots,0)\in\mathbb{R}^{N}$.
  5. 5.
    (Symmetries) If $\Omega$ is additionally radially symmetric, then $\phi_{1}$ is radially symmetric and $\phi_{2}$ is foliated Schwarz symmetric.
Part 1 is an analogue of the famous Courant nodal domain theorem [10]. Here, as usual, the connected components of $\mathbb{R}^{N}\smallsetminus\phi_{k}^{-1}(0)$ are called the nodal domains of $\phi_{k}$. We emphasize that the fact that each $\phi_{k}$ changes sign for every $k\geq 2$ is not an easy consequence of orthogonality (as in the bounded domain case with Dirichlet boundary conditions), because one must consider at all times the (sign-changing) weight $Q_{\Omega}$; instead, the proof is a consequence of Picone’s identity, see Lemma 2.7.
The proof of Part 2 is based on estimates for the Green function of the operator $\Delta+1$ in $\mathbb{R}^{N}$ adapting the ideas from [17] and on comparison principles. We mention that the same estimates for $\phi_{1}$ appeared in [15, equation (1.9)], where problem (1.2) is considered in the context of shape optimization.
Part 3 is a Faber-Krahn type inequality, this result is shown using symmetrizations and states that the first eigenvalue is minimized at balls (among open sets of fixed measure). In contrast, Part 4 proves that the second eigenvalue does not achieve its infimum among open bounded sets of fixed measure. This is sometimes referred to as the Hong–Krahn–Szegö inequality. Note that the inequality in (1.8) is strict, in contrast to what happens in the Dirichlet bounded domain case (where the equality is achieved).
Part 5 shows that the eigenfunctions inherit the symmetries of the domain. The radial symmetry of $\phi_{1}$ is an easy consequence of the simplicity of the first eigenvalue, whereas the foliated Schwarz symmetry of $\phi_{2}$ is shown using polarizations. Foliated Schwarz symmetry refers to an axially symmetric function together with some monotonicity properties in the polar angle; we refer to Section 2.2 for a precise definition.
To close this introduction, let us mention that we have only considered positive eigenvalues of (1.2). However, since the function $Q_{\Omega}$ is sign-changing one cannot rule out the existence of negative eigenvalues. The following is a nonexistence result of eigenfunctions in $H^{1}(\mathbb{R}^{N})$ associated to negative eigenvalues.
Theorem 1.7.
Let $\Omega\subset\mathbb{R}^{N}$ be a smooth, open and bounded domain and let $(\Lambda,u)$ be a nontrivial solution to (1.2). If $\Omega$ is starshaped (or the disjoint union of starshaped sets), then $\Lambda>0$.
The proof relies on a Pohožaev-type identity shown in [5]. The question remains open whether this is true for every $\Omega$ (not necessarily starshaped).
Question 1.8.
Is it true that, for any open bounded subset $\Omega\subset\mathbb{R}^{N}$, there are no solutions $(\Lambda,u)$ to problem (1.2) with $\Lambda<0$?
The paper is organized as follows. In Section 2 we study the existence of solutions to the eigenvalue problem (1.2) and show the qualitative properties stated in Theorem 1.6 and the nonexistence Theorem 1.7. Section 3 is devoted to the asymptotic analysis of the nonlinear problem as $p\to 2$. In Subsection 3.2, we prove our main result on uniqueness and nondegeneracy in the almost linear regime, that is, Theorem 1.1. In Section 4 we derive sharp decay estimates for the solutions to (1.1) for small values of $p$. Finally, Section 5 contains a multiplicity result of least energy solutions in the slightly subcritical regime.
2 The eigenvalue problem
In this section we study the eigenvalue problem, labelled (1.2) in the introduction,
$\begin{cases}-\Delta u=\Lambda Q_{\Omega}u,\\ u\in H^{1}(\mathbb{R}^{N}),\quad u\neq 0,\end{cases}$
where $\Omega$ is an open bounded subset of $\mathbb{R}^{N}$ and $Q_{\Omega}:\mathbb{R}^{N}\to\mathbb{R}$ is given by
$Q_{\Omega}(x)=\begin{cases}1&\text{if \ }x\in\Omega,\\ -1&\text{if \ }x\in\mathbb{R}^{N}\smallsetminus\Omega.\end{cases}$
A solution to this problem is a pair $(\Lambda,u)\in\mathbb{R}\times H^{1}(\mathbb{R}^{N})$ such that $u\neq 0$ and
$\int_{\mathbb{R}^{N}}\nabla u\cdot\nabla\varphi=\Lambda\int_{\mathbb{R}^{N}}Q_{\Omega}u\varphi,\qquad\text{for all \ }\varphi\in C^{\infty}_{c}(\mathbb{R}^{N}).$
2.1 Existence of positive eigenvalues
Let $\Lambda>0$. Since $(\Lambda,tu)$ with $t\in\mathbb{R}\smallsetminus\{0\}$ solves (1.2) whenever $(\Lambda,u)$ does, it suffices to look for solutions with $u$ belonging to the following set
$\mathcal{S}_{\Omega}:=\Big\{v\in D^{1,2}(\mathbb{R}^{N}):\int_{\mathbb{R}^{N}}Q_{\Omega}v^{2}=1\Big\}.$ (2.1)
Note that, if $v\in\mathcal{S}_{\Omega}$, then $v\in H^{1}(\mathbb{R}^{N})$, because $D^{1,2}(\mathbb{R}^{N})\subset L^{2}_{loc}(\mathbb{R}^{N})$ and, as $\Omega$ is bounded,
$\int_{\mathbb{R}^{N}\smallsetminus\Omega}v^{2}=\int_{\Omega}v^{2}-1<\infty.$ (2.2)
We write
$\langle u,v\rangle:=\int_{\mathbb{R}^{N}}\nabla u\cdot\nabla v\qquad\text{and}\qquad\|u\|^{2}:=\int_{\mathbb{R}^{N}}|\nabla u|^{2}$
for the inner product and norm in $D^{1,2}(\mathbb{R}^{N})$.
Lemma 2.1.
Let $V$ be a closed linear subspace of $D^{1,2}(\mathbb{R}^{N})$ such that $V\cap\mathcal{S}_{\Omega}\neq\emptyset$. Then,
$\Lambda_{V}:=\inf_{v\in V\cap\mathcal{S}_{\Omega}}\|v\|^{2}$
is attained at some function $\phi_{V}\in V\cap\mathcal{S}_{\Omega}$. As a consequence, $\phi_{V}\in H^{1}(\mathbb{R}^{N})$, and it satisfies
$\langle\phi_{V},v\rangle=\Lambda_{V}\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{V}v,\qquad\text{for all \ }v\in V\cap H^{1}(\mathbb{R}^{N}).$ (2.3)
Proof.
Let $u_{n}\in V\cap\mathcal{S}_{\Omega}$ be such that $\|u_{n}\|^{2}\to\Lambda_{V}$. Then $(u_{n})$ is bounded in $D^{1,2}(\mathbb{R}^{N})$ and, passing to a subsequence, $u_{n}\rightharpoonup u$ weakly in $D^{1,2}(\mathbb{R}^{N})$, $u_{n}\to u$ in $L^{2}_{loc}(\mathbb{R}^{N})$ and $u_{n}\to u$ a.e. in $\mathbb{R}^{N}$. As $V$ is a closed linear subspace of $D^{1,2}(\mathbb{R}^{N})$ we have that $u\in V$. Furthermore, as $u_{n}\in\mathcal{S}_{\Omega}$, using Fatou’s lemma we obtain
$1=\lim_{n\to\infty}\int_{\Omega}u_{n}^{2}-\lim_{n\to\infty}\int_{\mathbb{R}^{N}\smallsetminus\Omega}u_{n}^{2}\leq\int_{\Omega}u^{2}-\int_{\mathbb{R}^{N}\smallsetminus\Omega}u^{2}=\int_{\mathbb{R}^{N}}Q_{\Omega}u^{2}.$ (2.4)
Therefore, $u\neq 0$ and, as a consequence,
$0<\|u\|^{2}\leq\lim_{n\to\infty}\|u_{n}\|^{2}=\Lambda_{V}\leq\left\|\frac{u}{\sqrt{\int_{\mathbb{R}^{N}}Q_{\Omega}u^{2}}}\right\|^{2}=\frac{\|u\|^{2}}{\int_{\mathbb{R}^{N}}Q_{\Omega}u^{2}}.$
This implies that $\int_{\mathbb{R}^{N}}Q_{\Omega}u^{2}\leq 1$ and, together with (2.4) yields $u\in V\cap\mathcal{S}_{\Omega}$ and $\|u\|^{2}=\Lambda_{V}$. As shown in (2.2), $u\in H^{1}(\mathbb{R}^{N})$.
To prove that $\phi_{V}:=u$ satisfies (2.3) we fix $v\in V\cap H^{1}(\mathbb{R}^{N})$. Since $\{w\in H^{1}(\mathbb{R}^{N}):\int_{\mathbb{R}^{N}}Q_{\Omega}w^{2}>0\}$ is open in $H^{1}(\mathbb{R}^{N})$, there exists $\varepsilon>0$ such that $\int_{\mathbb{R}^{N}}Q_{\Omega}(u+tv)^{2}>0$ for all $t\in(-\varepsilon,\varepsilon)$. Then,
$\frac{u+tv}{\sqrt{\int_{\mathbb{R}^{N}}Q_{\Omega}(u+tv)^{2}}}\in V\cap\mathcal{S}_{\Omega}\quad\text{for all \ }t\in(-\varepsilon,\varepsilon),$
and 0 is a minimum of the function
$f(t):=\left\|\frac{u+tv}{\sqrt{\int_{\mathbb{R}^{N}}Q_{\Omega}(u+tv)^{2}}}\right\|^{2}=\frac{\|u+tv\|^{2}}{\int_{\mathbb{R}^{N}}Q_{\Omega}(u+tv)^{2}},\qquad t\in(-\varepsilon,\varepsilon).$
This function is differentiable and
$0=f^{\prime}(0)=2\left(\langle u,v\rangle-\Lambda_{V}\int_{\mathbb{R}^{N}}Q_{\Omega}uv\right).$
This completes the proof. ∎
Lemma 2.2.
If $\{\phi_{k}\in\mathcal{S}_{\Omega}:k\in\mathbb{N}\}$ is a set of functions that are orthogonal in $D^{1,2}(\mathbb{R}^{N})$, $\Lambda_{k}:=\|\phi_{k}\|^{2}$ and $(\Lambda_{k},\phi_{k})$ solves (1.2), then $(\Lambda_{k})$ is unbounded. As a consequence, for each $\Lambda>0$, the space
[…]
has finite dimension.
Proof.
As $\{\Lambda_{k}^{-\frac{1}{2}}\phi_{k}:k\in\mathbb{N}\}$ is an orthonormal set in $D^{1,2}(\mathbb{R}^{N})$, passing to a subsequence we have that $\Lambda_{k}^{-\frac{1}{2}}\phi_{k}\rightharpoonup 0$ weakly in $D^{1,2}(\mathbb{R}^{N})$ and $\Lambda_{k}^{-\frac{1}{2}}\phi_{k}\to 0$ strongly in $L^{2}_{loc}(\mathbb{R}^{N})$. Therefore,
$0\leq\Lambda_{k}^{-1}=\Lambda_{k}^{-2}\|\phi_{k}\|^{2}=\Lambda_{k}^{-1}\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{k}^{2}\leq\int_{\Omega}(\Lambda_{k}^{-\frac{1}{2}}\phi_{k})^{2}\longrightarrow 0.$
This proves that $(\Lambda_{k})$ is unbounded.
If $E_{\Lambda}$ had infinite dimension, it would contain a set $\{\phi_{k}\in\mathcal{S}_{\Omega}:k\in\mathbb{N}\}$ of orthogonal functions in $D^{1,2}(\mathbb{R}^{N})$ with $\Lambda=\|\phi_{k}\|^{2}$. Arguing as before, this would imply that $\Lambda=0$, a contradiction. ∎
We are ready to prove Theorem 1.3.
Proof of Theorem 1.3.
We apply Lemma 2.1 inductively to the spaces $V_{k}$, where $V_{1}:=D^{1,2}(\mathbb{R}^{N})$ and $V_{k}$ is the orthogonal complement in $D^{1,2}(\mathbb{R}^{N})$ of the space generated by $\{\phi_{1},\ldots,\phi_{k-1}\}$ if $k\geq 2$. Lemma 2.1 states that there exists $\phi_{k}\in V_{k}\cap\mathcal{S}_{\Omega}$ such that $\|\phi_{k}\|^{2}=\Lambda_{k}$ and
$\langle\phi_{k},v\rangle=\Lambda_{k}\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{k}v\qquad\text{for all \ }v\in V_{k}\cap H^{1}(\mathbb{R}^{N}).$ (2.5)
Then, $\langle\phi_{k},\phi_{i}\rangle=0$ for all $i=1,\ldots,k-1$ and, as $0=\langle\phi_{i},\phi_{k}\rangle=\Lambda_{i}\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{i}\phi_{k}$ and $\Lambda_{i}>0$ we get that
$\langle\phi_{k},\phi_{i}\rangle=0=\Lambda_{k}\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{k}\phi_{i}\qquad\text{for all \ }i=1,\ldots,k-1.$ (2.6)
$(i):$ (2.5) and (2.6) imply that $(\Lambda_{k},\phi_{k})$ solves (1.2).
$(ii):$ This follows from standard regularity arguments.
$(iii):$ Replacing $\phi_{1}$ with $\phi_{1}$ if necessary, we may assume that $\phi_{1}^{+}:=\max\{\phi_{1},0\}\neq 0$. Testing equation (1.2) with $\phi_{1}^{+}$ we get
$0<\|\phi_{1}^{+}\|^{2}=\Lambda_{1}\int_{\mathbb{R}^{N}}Q_{\Omega}(\phi_{1}^{+})^{2}.$
Then,
$\widehat{\phi}_{1}^{+}:=\frac{\phi_{1}^{+}}{\sqrt{\int_{\mathbb{R}^{N}}Q_{\Omega}(\phi_{1}^{+})^{2}}}\in\mathcal{S}_{\Omega}\qquad\text{and}\qquad\Lambda_{1}=\inf_{v\in\mathcal{S}_{\Omega}}\|v\|^{2}=\|\widehat{\phi}_{1}^{+}\|^{2}.$
It follows from Lemma 2.1 that $(\Lambda_{1},\widehat{\phi}_{1}^{+})$ solves (1.2). The strong maximum principle for strong solutions [18, Theorem 9.6] applied to the equation
$\Delta\widehat{\phi}_{1}^{+}+\Lambda_{1}\mathds{1}_{\mathbb{R}^{N}\smallsetminus\Omega}\widehat{\phi}_{1}^{+}=\Lambda_{1}\mathds{1}_{\Omega}\widehat{\phi}_{1}^{+}\geq 0\qquad\text{in \ }\mathbb{R}^{N},$
where $\mathds{1}_{A}$ is the characteristic function of $A$, yields $\widehat{\phi}_{1}^{+}>0$ in $\mathbb{R}^{N}$. Hence, $\phi_{1}=\phi_{1}^{+}>0$ in $\mathbb{R}^{N}$.
$(iv):$ As $V_{1}\supset V_{2}\supset\cdots\supset V_{k}\supset\cdots$ we have that $0<\Lambda_{1}\leq\Lambda_{2}\leq\cdots\leq\Lambda_{k}\leq\cdots$ and, by Lemma 2.2, $\Lambda_{k}\to\infty$ and $\dim(E_{k})<\infty$. If $\Lambda_{1}=\Lambda_{2}$ there would exist two functions $v_{1},v_{2}\in\mathcal{S}_{\Omega}$, orthogonal in $L^{2}(\mathbb{R}^{N})$, such that $(\Lambda_{1},v_{1})$ and $(\Lambda_{1},v_{2})$ solve (1.2). But then, by $(iii)$, their product $v_{1}v_{2}$ would be either strictly positive or strictly negative in $\mathbb{R}^{N}$, contradicting $\int_{\mathbb{R}^{N}}v_{1}v_{2}=0$. ∎
Remark 2.3.
The set of eigenfunctions $\{\phi_{k}\in\mathcal{S}_{\Omega}:\;k\in\mathbb{N}\,\}$ does not generate the space $D^{1,2}(\mathbb{R}^{N})$. Indeed, let $\varphi\in C_{c}^{\infty}(\mathbb{R}^{N})$ be such that $\int_{\mathbb{R}^{N}}Q_{\Omega}\varphi^{2}=-1$ and suppose, by contradiction, that $\varphi_{m}:=\sum_{i=1}^{m}a_{i}\phi_{i}\to\varphi$ in $D^{1,2}(\mathbb{R}^{N})$ for some $a_{i}\in\mathbb{R}$. Passing to a subsequence, $\varphi_{m}\to\varphi$ strongly in $L^{2}(\Omega)$ and a.e. in $\mathbb{R}^{N}$. Now, since $\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{i}\phi_{j}=0$ if $i\neq j$ and $\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{i}^{2}=1$,
$\int_{\mathbb{R}^{N}}Q_{\Omega}\varphi_{m}^{2}=\sum_{i=1}^{m}a_{i}^{2}\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{i}^{2}=\sum_{i=1}^{m}a_{i}^{2}\geq 0\qquad\text{ for every }m\in\mathbb{N}.$
Then, using Fatou’s Lemma,
$0\leq\liminf_{m\to\infty}\int_{\mathbb{R}^{N}}Q_{\Omega}\varphi_{m}^{2}\leq\int_{\Omega}\varphi^{2}-\int_{\mathbb{R}^{N}\setminus\Omega}\varphi^{2}=\int_{\mathbb{R}^{N}}Q_{\Omega}\varphi^{2}=-1,$
a contradiction.
2.2 Properties of the eigenfunctions
We show next that the eigenfunctions $\phi_{k}$ given by Theorem 1.3 share some of the well known properties of Dirichlet eigenfunctions in bounded domains. Recall that $\phi_{k}$ denotes the $k$-th eigenfunction associated to the $k$-th eigenvalue $\Lambda_{k}$ normalized so that
$\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{k}^{2}=1.$ (2.7)
2.2.1 Monotonicity and scalings
We begin with some easy-to-check properties regarding scalings and monotonicity.
Remark 2.4.
The following properties follow directly from the definition and a change of variables.
  1. 1.
    If $\Omega_{1},\Omega_{2}\subset\mathbb{R}^{N}$ are open bounded subsets and $\Omega_{1}\subset\Omega_{2}$, then $\Lambda_{1}(\Omega_{1})\geq\Lambda_{1}(\Omega_{2})$.
  2. 2.
    If $t>0$, then $\Lambda_{k}(t\Omega)=t^{-2}\Lambda_{k}(\Omega)$.
2.2.2 Decay
Next, we show the sharp decay of $\phi_{k}$.
Theorem 2.5.
For each $k\in\mathbb{N}$ there exists $C>0$ such that
$|\phi_{k}|\leq C|x|^{-\frac{N-1}{2}}e^{-\sqrt{\Lambda_{k}}|x|}\qquad\text{for all \ }|x|\geq 1.$ (2.8)
This estimate is sharp, in the sense that there exists $c>0$ such that
$|\phi_{1}|>c|x|^{-\frac{N-1}{2}}e^{-\sqrt{\Lambda_{1}}|x|}\qquad\text{for all \ }|x|\geq 1.$
Proof.
Let $R>0$ be such that $\Omega\subset B_{R}(0)$. Then,
$\Delta\phi_{k}(x)=-\Lambda_{k}Q_{\Omega}(x)\phi_{k}(x)=\Lambda_{k}\phi_{k}(x)\quad\text{for all \ }x\in\mathbb{R}^{N}\smallsetminus B_{R}(0).$
Setting $h(r):=r^{-\frac{N-1}{2}}e^{-\sqrt{\Lambda_{k}}r}$ and $u(x):=h(|x|)$ for $|x|=r>0$, we see that
$-\Delta u(x)+\Lambda_{k}u(x)=-h^{\prime\prime}(r)-\frac{N-1}{r}h^{\prime}(r)+\Lambda_{k}h(r)=\frac{(N-3)(N-1)}{4}e^{-\sqrt{\Lambda_{k}}r}r^{-\frac{N+3}{2}}\geq 0.$
Fix $a>0$ such that $w(x):=au(x)\pm\phi_{k}(x)>0$ on $\partial B_{R}(0).$ As $\Delta w+\Lambda_{k}w\geq 0$ in $\mathbb{R}^{N}\smallsetminus B_{R}(0)$, we have, by the maximum principle, that $w>0$ in $\mathbb{R}^{N}\smallsetminus B_{R}(0)$, which yields (2.8).
For the sharp estimate, assume first that $\Lambda_{1}=1.$ The Green function for the operator $\Delta+1$ in $\mathbb{R}^{N}$ is given by $G(x,y)=|x-y|^{-\frac{N-2}{2}}K_{\frac{N-2}{2}}(|x-y|)$, where $K_{\nu}$ is the modified Bessel function of order $\nu$; see [17, Section 4]. Since $\Delta\phi_{1}+\phi_{1}=2\phi_{1}\mathds{1}_{\Omega}$ in $\mathbb{R}^{N}$, where $\mathds{1}_{\Omega}$ is the characteristic function of $\Omega$, we have that
$\phi_{1}(x)=2\int_{\Omega}G(x,y)\phi_{1}(y)\,\mathrm{d}y\geq 2|\Omega|\left(\inf_{\Omega}\phi_{1}\right)\inf_{y\in\Omega}G(x,y).$
The claim for $\Lambda_{1}=1$ now follows from the asymptotic expansions for $K_{\frac{N-2}{2}}$ given in [17, Appendix C]. The estimate for general $\Lambda_{1}>0$ follows by rescaling. ∎
2.2.3 Nodal domains
Next we give a bound for the number of nodal domains of $\phi_{k}$ (the connected components of $\mathbb{R}^{N}\smallsetminus\phi_{k}^{-1}(0)$).
Proposition 2.6.
$\phi_{k}$ has at most $k$ nodal domains.
Proof.
Assume that $\phi_{k}$ has $m$ nodal domains and that $m\geq k\geq 2$. Choose $k$ nodal domains $U_{1},\ldots,U_{k}$ and define
$w_{i}(x):=\begin{cases}\phi_{k}(x)&\text{if }x\in U_{i},\\ 0&\text{if }x\in\mathbb{R}^{N}\smallsetminus U_{i}.\end{cases}$
Then, $w_{i}\in D^{1,2}(\mathbb{R}^{N})$,
$0<\|w_{i}\|^{2}=\langle\phi_{k},w_{i}\rangle=\Lambda_{k}\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{k}w_{i}=\Lambda_{k}\int_{\mathbb{R}^{N}}Q_{\Omega}w_{i}^{2}\qquad\text{for each \ }i=1,\ldots,k,$ (2.9)
and $\langle w_{i},w_{j}\rangle=0$ if $i\neq j$. Hence, the linear subspace generated by $w_{1},\ldots,w_{k}$ has dimension $k$ and, as a consequence, there exists $w=c_{1}w_{1}+\cdots+c_{k}w_{k}\neq 0$, $c_{i}\in\mathbb{R}$, such that
$\langle w,\phi_{j}\rangle=0\qquad\text{for every \ }j=1,\ldots,k-1.$
It follows from (2.9) that
$\int_{\mathbb{R}^{N}}Q_{\Omega}w^{2}=\sum_{i=1}^{k}\int_{U_{i}}Q_{\Omega}w^{2}=\sum_{i=1}^{k}c_{i}^{2}\int_{\mathbb{R}^{N}}Q_{\Omega}w^{2}_{i}>0.$
Setting $t:=\Big(\int_{\mathbb{R}^{N}}Q_{\Omega}w^{2}\Big)^{-1/2}$ and $u:=tw$ we have that
$u\in V_{k}\cap\mathcal{S}_{\Omega},$
where $V_{k}$ is the orthogonal complement in $D^{1,2}(\mathbb{R}^{N})$ of $\mathrm{span}\{\phi_{1},\ldots,\phi_{k-1}\}$, and from (2.9) we get
$\|u\|^{2}=\sum_{i=1}^{k}\|tc_{i}w_{i}\|^{2}=\sum_{i=1}^{k}\Lambda_{k}\int_{\mathbb{R}^{N}}Q_{\Omega}(tc_{i}w_{i})^{2}=\Lambda_{k}\int_{\mathbb{R}^{N}}Q_{\Omega}u^{2}.$
This implies that $u$ solves the eigenvalue problem (1.2) with $\Lambda=\Lambda_{k}$.
If $m>k$, then $A:=\mathbb{R}^{N}\smallsetminus(U_{1}\cup\cdots\cup U_{k})$ has nonempty interior. As $u$ vanishes in $A$, the unique continuation principle [16, Theorem 1.2] leads to a contradiction. Therefore, $m\leq k$, as claimed. ∎
Lemma 2.7.
$\phi_{k}$ changes sign for every $k\geq 2$.
Proof.
We use Picone’s identity which says that, for any $u,v\in C^{1}(\mathbb{R}^{N})$ with $v>0$ and $u\geq 0$,
$|\nabla u|^{2}-\nabla\left(\frac{u^{2}}{v}\right)\cdot\nabla v=|\nabla u|^{2}+\frac{u^{2}}{v^{2}}|\nabla v|^{2}-2\frac{u}{v}\nabla u\cdot\nabla v\geq 0,$ (2.10)
(see, for instance, [2, Ec. (1)]).
Assume, by contradiction, that $\phi_{k}\geq 0$ for some $k\geq 2$. Since $\Delta\phi_{k}+\Lambda_{k}\mathds{1}_{\mathbb{R}^{N}\smallsetminus\Omega}\phi_{k}=\Lambda_{k}\mathds{1}_{\Omega}\phi_{k}\geq 0$ in $\mathbb{R}^{N}$, the strong maximum principle [18, Theorem 9.6] yields that $\phi_{k}>0$ in $\mathbb{R}^{N}$. For $\varepsilon>0$, let $\psi_{\varepsilon}=\frac{\phi_{1}^{2}}{\phi_{k}+\varepsilon}$. Note that $\psi_{\varepsilon}\in H^{1}(\mathbb{R}^{N})$. Indeed, by Theorem 1.3, $\psi_{\varepsilon}\in C^{1}(\mathbb{R}^{N})$. Moreover, $0<\psi_{\varepsilon}<\frac{\phi_{1}^{2}}{\varepsilon}$ and
$|\nabla\psi_{\varepsilon}|=\left|\frac{2\phi_{1}\nabla\phi_{1}}{\phi_{k}+\varepsilon}-\frac{\phi_{1}^{2}\nabla\phi_{k}}{(\phi_{k}+\varepsilon)^{2}}\right|\leq 2\varepsilon^{-1}\phi_{1}|\nabla\phi_{1}|+\varepsilon^{-2}\phi_{1}^{2}|\nabla\phi_{k}|\in L^{2}(\mathbb{R}^{N}),$
by Theorem 2.5 and Hölder’s inequality. Thus, testing with $\psi_{\epsilon}$ and using (2.10),
$\Lambda_{k}\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{k}\psi_{\varepsilon}=\int_{\mathbb{R}^{N}}\nabla\phi_{k}\cdot\nabla\psi_{\varepsilon}=\int_{\mathbb{R}^{N}}\nabla(\phi_{k}+\epsilon)\cdot\nabla\left(\frac{\phi_{1}^{2}}{\phi_{k}+\epsilon}\right)\leq\int_{\mathbb{R}^{N}}|\nabla\phi_{1}|^{2}=\Lambda_{1}.$
Since $\psi_{\varepsilon}\phi_{k}=\frac{\phi_{k}}{\phi_{k}+\epsilon}\phi_{1}^{2}\to\phi_{1}^{2}$ pointwisely as $\varepsilon\to 0$ and $\frac{\phi_{k}}{\phi_{k}+\epsilon}\phi_{1}^{2}\leq\phi_{1}^{2}$ for every $\varepsilon>0$, dominated convergence implies that $\Lambda_{k}\leq\Lambda_{1}$, which contradicts Theorem 1.3. ∎
2.2.4 Eigenfunctions, eigenvalues, and symmetries
We now prove the Faber-Krahn type inequality stated in the introduction.
Theorem 2.8 (Faber-Krahn inequality).
For every $\Omega\subset\mathbb{R}^{N}$ open bounded set, it holds that
$\Lambda_{1}(\Omega)\geq\Lambda_{1}(\Omega^{*}),$ (2.11)
where $\Omega^{*}$ is the ball centered at the origin with the same volume as $\Omega$. Moreover, this inequality is an equality if and only if $\Omega$ is a ball.
Proof.
Let $\phi_{1}$ be the positive eigenfunction with eigenvalue $\Lambda_{1}$ for the problem (1.2) in $\Omega$ that satisfies
$\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{2}=1.$
Let $\phi_{1}^{*}$ be the Schwarz symmetrization of $\phi_{1}$ (also called symmetric-decreasing rearrangement). As shown in [19, Section 3.3],
$\int_{\mathbb{R}^{N}}\phi_{1}^{2}=\int_{\mathbb{R}^{N}}(\phi_{1}^{*})^{2},$
$(\mathds{1}_{\Omega})^{*}=\mathds{1}_{\Omega^{*}}$ and $(\phi_{1}^{2})^{*}=(\phi_{1}^{*})^{2}$. Then, using [19, Theorem 3.4] we see that
$\int_{\Omega}\phi_{1}^{2}\leq\int_{\Omega^{*}}(\phi_{1}^{*})^{2}.$
Therefore,
$1=\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{2}=2\int_{\Omega}\phi_{1}^{2}-\int_{\mathbb{R}^{N}}\phi_{1}^{2}\leq 2\int_{\Omega^{*}}(\phi_{1}^{*})^{2}-\int_{\mathbb{R}^{N}}(\phi_{1}^{*})^{2}=\int_{\mathbb{R}^{N}}Q_{\Omega^{*}}(\phi_{1}^{*})^{2}.$
By [19, Lemma 7.17] we have that
$\|\nabla\phi_{1}^{\ast}\|\leq\|\nabla\phi_{1}\|.$
As a consequence,
$\Lambda_{1}=\int_{\mathbb{R}^{N}}|\nabla\phi_{1}|^{2}\geq\int_{\mathbb{R}^{N}}|\nabla\phi_{1}^{*}|^{2}\geq\frac{\int_{\mathbb{R}^{N}}|\nabla\phi_{1}^{\ast}|^{2}}{\int_{\mathbb{R}^{N}}Q_{\Omega^{*}}(\phi_{1}^{*})^{2}}\geq\Lambda_{1}^{\ast},$
as claimed.
If $\Lambda_{1}=\Lambda_{1}^{*}$, then $\|\nabla\phi_{1}^{\ast}\|=\|\nabla\phi_{1}\|$. As shown in [4, Theorem 1.1], this implies that $\phi_{1}$ is a translation of $\phi_{1}^{*}$, i.e., $\phi_{1}^{*}=\phi_{1}(\,\cdot\,-\zeta)$ for some $\zeta\in\mathbb{R}^{N}$. Furthermore, a change of variable gives
$\Lambda_{1}\int_{\mathbb{R}^{N}}Q_{\Omega+\zeta}(\phi_{1}^{*})^{2}=\Lambda_{1}\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{2}=\|\nabla\phi_{1}\|^{2}=\|\nabla\phi_{1}^{\ast}\|^{2}=\Lambda_{1}^{*}\int_{\mathbb{R}^{N}}Q_{\Omega^{*}}(\phi_{1}^{*})^{2}=\Lambda_{1}\int_{\mathbb{R}^{N}}Q_{\Omega^{*}}(\phi_{1}^{*})^{2}.$
This implies that $Q_{\Omega+\zeta}=Q_{\Omega^{*}}$. Hence, $\Omega+\zeta=\Omega^{*}$, i.e., $\Omega$ is a ball, as claimed. ∎
If $\Omega$ is radially symmetric, the first and second eigenfunctions have the expected symmetries.
Proposition 2.9.
If $\Omega$ is a bounded open radially symmetric set, then $\phi_{1}$ is radially symmetric.
Proof.
For any reflection $\varrho$ on a hyperplane $V$ through the origin, the pair $(\Lambda_{1},\phi_{1}\circ\varrho)$ is a solution to (1.2). Since $\dim(E_{1})=1$, $\phi_{1}(\varrho x)=\phi_{1}(x)$ for all $x\in V$ and $|\phi_{1}|>0$ in $\mathbb{R}^{N}$, we have that $\phi_{1}\circ\varrho=\phi_{1}$. As a consequence, since for any $x,y\in\mathbb{R}^{N}$ such that $x\neq y$ and $|x|=|y|\neq 0$, the reflection on the hyperplane through the origin orthogonal to xyx-y maps $x$ to $y$, we have that $\phi_{1}(x)=\phi_{1}(y)$. This proves that $\phi_{1}$ is radial. ∎
Next, we show that the eigenfunctions with eigenvalue $\Lambda_{2}$ are foliated Schwarz symmetric. Let us recall what this means.
Let $\mathbb{S}^{N-1}=\{x\in\mathbb{R}^{N}\,:\,|x|=1\}$ be the unit sphere. For each $e\in\mathbb{S}^{N-1}$ we consider the half-space $H(e):=\{x\in\mathbb{R}^{N}\,:\,x\cdot e>0\}$, and we denote by $\phi_{e}$ the composition of the function $\phi:\mathbb{R}^{N}\to\mathbb{R}$ with the reflection $\sigma_{e}$ with respect to the hyperplane $\partial H(e)$, that is
$\phi_{e}:\mathbb{R}^{N}\to\mathbb{R}\quad\text{is given by}\quad\phi_{e}(x):=\phi(\sigma_{e}(x)),\quad\text{where}\;\;\;\sigma_{e}(x)=x-2(x\cdot e)e.$
The polarization $\phi_{H}:\mathbb{R}^{N}\to\mathbb{R}$ with respect to the half-space $H:=H(e)$ is defined by
$\phi_{H}:=\begin{cases}\max\{\phi,\phi_{e}\}&\text{in}\;\;\overline{H},\\ \min\{\phi,\phi_{e}\}&\text{in}\;\;\mathbb{R}^{N}\smallsetminus H.\end{cases}$
A function $u\in C^{0}(\Omega)$ is said to be foliated Schwarz symmetric with respect to $e^{\ast}\in\mathbb{S}^{N-1}$ if $u$ is axially symmetric with respect to the axis $\mathbb{R}e^{\ast}:=\{te^{\ast}\,:\,t\in\mathbb{R}\}$ (i.e., it is invariant under rotations around this axis), and nonincreasing in the polar angle $\theta:=\arccos(\frac{x}{|x|}\cdot e^{\ast})\in\,[0,\pi]$.
We use the following characterization of foliated Schwarz symmetry.
Lemma 2.10.
Let $\Omega\subset\mathbb{R}^{N}$ be a radially symmetric open set. There exists $e^{*}\in\mathbb{S}^{N-1}$ such that $u\in C^{0}(\overline{\Omega})$ is foliated Schwarz symmetric with respect to $e^{*}$ if and only if for every $e\in\mathbb{S}^{N-1}$ either
[…]
Proof.
The proof is given in [3] (see also [22, Proposition 3.2] and [26]). This proof only considers the case where $\Omega$ is connected, but the general case follows easily since each component satisfies the same inequalities that completely determine the axis of symmetry and angular monotonicity. ∎
We also use the following simple fact.
Lemma 2.11.
Let $\varphi$ and $f$ Lebesgue-measurable functions in $\mathbb{R}^{N}$ such that $\varphi$ is radially symmetric and $\int_{\mathbb{R}^{N}}|\varphi f|<\infty$, then $\int_{\mathbb{R}^{N}}\varphi f_{H}=\int_{\mathbb{R}^{N}}\varphi f.$
Proof.
Let $M=\{x\in\mathbb{R}^{N}\,:\;f(x)=f_{H}(x)\}$. Notice that $\sigma_{e}(M)=M$ and $f_{H}=f\circ\sigma_{e}$ in $M^{c}$, then
$\int_{\mathbb{R}^{N}}\varphi f_{H}=\int_{M}\varphi f+\int_{\mathbb{R}^{N}\smallsetminus M}\varphi(f\circ\sigma_{e})=\int_{\mathbb{R}^{N}}\varphi f$
as claimed. ∎
The following result describes the symmetry of the second eigenfunctions.
Proposition 2.12.
Let $\Omega$ be a radially symmetric open bounded subset of $\mathbb{R}^{N}$. Then any eigenfunction of (1.2) with eigenvalue $\Lambda_{2}$ is foliated Schwarz symmetric.
Proof.
Let $\phi_{1}$ and $\phi$ be eigenfunctions with eigenvalues $\Lambda_{1}$ and $\Lambda_{2}$, respectively, such that $\phi_{1},\phi\in\mathcal{S}_{\Omega}$. Notice that $\Delta\phi_{1}=\Lambda_{1}Q_{\Omega}\phi_{1}$ is radially symmetric. By Lemma 2.11,
$\int_{\mathbb{R}^{N}}\nabla\phi_{1}\cdot\nabla\phi_{H}=\int_{\mathbb{R}^{N}}-(\Delta\phi_{1})\phi_{H}=\int_{\mathbb{R}^{N}}-(\Delta\phi_{1})\phi=\int_{\mathbb{R}^{N}}\nabla\phi_{1}\cdot\nabla\phi=0.$
Therefore, $\phi_{H}$ is an eigenfunction of (1.2) with eigenvalue $\Lambda_{2}$. Notice that $|\phi-\phi_{e}|=2\phi_{H}-(\phi+\phi_{e})$ in $H:=H(e)$ and $|\phi-\phi_{e}|=2\phi_{H}-(\phi+\phi_{e})$ in $\mathbb{R}^{N}\smallsetminus H(e)$. By standard elliptic regularity arguments, $\phi,\phi_{H}\in W_{loc}^{2,s}(\mathbb{R}^{N})\cap C_{loc}^{1,\alpha}(\mathbb{R}^{N})$ for all $s\in[1,\infty)$ and $\alpha\in(0,1)$. Note that $2\phi_{H}-(\phi+\phi_{e})=0=-(2\phi_{H}-(\phi+\phi_{e}))$ on $\partial H(e)$, $|\phi-\phi_{e}|\in H^{1}(\mathbb{R}^{N})\cap W_{loc}^{1,s}(\mathbb{R}^{N})$ for all $s\geq 1$, and
$-\Delta|\phi-\phi_{e}|$ $=\begin{cases}\Lambda_{2}Q_{\Omega}[(\phi_{H}-\phi)+(\phi_{H}-\phi_{e})]&\text{in }\;H(e),\\ \Lambda_{2}Q_{\Omega}[(\phi-\phi_{H})+(\phi_{e}-\phi_{H})]&\text{in}\;\mathbb{R}^{N}\smallsetminus H(e),\end{cases}$
[…]
$=\Lambda_{2}Q_{\Omega}(\mathds{1}_{H(e)}-\mathds{1}_{\mathbb{R}^{N}\smallsetminus H(e)})|\phi-\phi_{e}|$ (2.12)
Since $|\Lambda_{2}Q_{\Omega}(\mathds{1}_{H(e)}-\mathds{1}_{\mathbb{R}^{N}\setminus H(e)})|=\Lambda_{2}\in L^{\infty}(\mathbb{R}^{N})$, standard elliptic regularity yields that
$w_{e}:=|\phi-\phi_{e}|\in W^{2,s}_{loc}(\mathbb{R}^{N})\cap C_{loc}^{1,\alpha}(\mathbb{R}^{N})\cap C^{\infty,\alpha}(H(e)\cap\Omega)\cap C_{loc}^{\infty,\alpha}(H(e)\cap(\mathbb{R}^{N}\smallsetminus\Omega))$ (2.13)
for all $s\geq 1,\ \alpha\in(0,1)$. Hence, $w_{e}$ is a strong solution to (2.2.4); in particular,
$-\Delta w_{e}\geq 0\quad\text{in $H(e)\cap\Omega$}\qquad\text{ and }\qquad-\Delta w_{e}+\Lambda_{2}w_{e}=0\quad\text{ in $H(e)\cap(\mathbb{R}^{N}\smallsetminus\Omega)$.}$
By the strong maximum principle, either $w_{e}\equiv 0$ or $w_{e}>0$ in each connected component of $H(e)\cap\Omega$. We claim that, if $w_{e}>0$ in some connected component $A\subset\Omega$, then $w_{e}>0$ in $H(e)$. Indeed, let $B$ be the connected component of the positive nodal set $\{x\in\mathbb{R}^{n}\>:\>w_{e}>0\}$ which contains $A$ and assume, by contradiction, that there exists $x_{0}\in\mathbb{R}^{N}$ such that $w_{e}(x_{0})=0$. Then there exists $y_{0}\in\partial B$ and a small ball $U$ with $y_{0}\in\partial U$, which is contained either in $B\cap\Omega$ or in $B\cap\mathbb{R}^{N}\smallsetminus\Omega$. In any case, by the Hopf Lemma, $\nabla w_{e}(y_{0})\neq 0.$ However, making a Taylor expansion around $y_{0}$, in view of (2.13), it follows that
$0\leq|\phi-\phi_{e}|=w_{e}(x)=\nabla w_{e}(y_{0})\cdot(x-y_{0})+o(|x-y_{0}|)\mbox{ as }|x-y_{0}|\to 0.$
But this implies that $\nabla w_{e}(y_{0})=0$, which yields a contradiction. Thus, $w_{e}>0$ in $H(e)$ as claimed. This shows that either $\phi\geq\phi_{e}$ or $\phi\leq\phi_{e}$ in $H(e)$ and the claim follows from Lemma 2.10. ∎
Recall that $\phi_{2}$ denotes an eigenfunction related to the second eigenvalue $\Lambda_{2}:=\Lambda_{2}(\Omega)$ such that $\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{2}=1$. Let
$A_{+}:=\{x\in\mathbb{R}^{N}\,:\,\phi_{2}(x)>0\}\qquad\text{and}\qquad A_{-}:=\{x\in\mathbb{R}^{N}\,:\,\phi_{2}(x)<0\}.$
By Lemma 2.7, $A_{\pm}\neq\emptyset$. The following result is one of the main ingredients for our Hong–Krahn–Szegö-type inequality.
Lemma 2.13.
Let $\Omega_{\pm}:=\Omega\cap A_{\pm}$ and let $\Lambda_{1}(\Omega_{+})$ and $\Lambda_{1}(\Omega_{-})$ denote the first eigenvalues associated to $\Omega_{+}$ and $\Omega_{-}$, respectively. Then,
$\Lambda_{2}(\Omega)>\max\{\Lambda_{1}(\Omega_{+}),\Lambda_{1}(\Omega_{-})\}.$ (2.14)
Proof.
We show first that $\Omega_{\pm}\neq\emptyset$. By Lemma 2.7, $\phi_{2}$ changes sign. Since $\lim_{|x|\to\infty}|\phi_{2}(x)|=0$, there are $a,b\in\mathbb{R}^{N}$ such that
$\phi_{2}(a)=\sup_{x\in\mathbb{R}^{N}}\phi_{2}(x)>0\quad\text{and}\quad\phi_{2}(b)=\inf_{x\in\mathbb{R}^{N}}\phi_{2}(x)<0.$
If $a\in\partial\Omega$, then $\Omega_{+}\neq\emptyset$ by continuity. If $a\in\mathbb{R}^{N}\setminus\partial\Omega$, then $\Lambda_{2}Q_{\Omega}(a)\phi_{2}(a)=-\Delta\phi_{2}(a)\geq 0$, hence $Q_{\Omega}(a)>0$ and, therefore, $a\in\Omega$, which also yields that $\Omega_{+}\neq\emptyset$. Arguing similarly we also have that $\Omega_{-}\neq\emptyset$.
Testing with $\phi_{2}^{+}=\max\{\phi_{2},0\}$,
$\|\phi_{2}^{+}\|^{2}=\Lambda_{2}\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{2}(\phi_{2}^{+})=\Lambda_{2}\int_{\mathbb{R}^{N}}(Q_{\Omega}\mathds{1}_{A_{+}})(\phi_{2}^{+})^{2}=\Lambda_{2}\int_{\mathbb{R}^{N}}Q_{\Omega_{+}}(\phi_{2}^{+})^{2},$
because $Q_{\Omega}\mathds{1}_{A_{+}}=\mathds{1}_{\Omega\cap A_{+}}-\mathds{1}_{(\mathbb{R}^{N}\setminus\Omega)\cap A_{+}}$ and $\phi_{2}^{+}=0$ in $\mathbb{R}^{N}\setminus A_{+}$. Then $\phi:=\phi_{2}^{+}/\sqrt{\int_{\mathbb{R}^{N}}Q_{\Omega_{+}}(\phi_{2}^{+})^{2}}\in S_{\Omega_{+}}$ (recall (2.1)) and
$\Lambda_{1}(\Omega_{+})=\inf_{v\in S_{\Omega_{+}}}\|v\|^{2}\leq\|\phi\|^{2}=\Lambda_{2}.$ (2.15)
Since $\phi\equiv 0$ in $A_{-}$, the inequality in (2.15) must be strict (by unique continuation). We obtain the same estimate for $\Lambda_{2}(\Omega_{-})$ using similar arguments and the claim follows. ∎
We denote by $B_{r}$ the open ball of radius $r$ centered at the origin in $\mathbb{R}^{N}$ and by
$r(c):=\left(\frac{c}{|B_{1}|}\right)^{1/N}$
the radius of the ball of volume $c$ in $\mathbb{R}^{N}$. As before,
$\Omega^{*}:=B_{r(|\Omega|)}.$
Theorem 2.14.
  • $(i)$
    For any open bounded subset $\Omega$ of $\mathbb{R}^{N}$,
    $\Lambda_{2}(\Omega)>\Lambda_{1}(B_{r(|\Omega|/2)}).$
  • $(ii)$
    For any $c>0$,
    $\inf\{\Lambda_{2}(\Omega):\Omega\subset\mathbb{R}^{N}\text{ is open and bounded and }|\Omega|=c\,\}=\Lambda_{1}(B_{r(c/2)}).$
    A minimizing sequence is given by $\Omega_{n}:=B_{r(c/2)}(-ne_{1})\cup B_{r(c/2)}(ne_{1})$, where $e_{1}=(1,0,\ldots,0)\in\mathbb{R}^{N}$.
Therefore, the second eigenvalue does not achieve its infimum among open bounded sets of fixed volume.
Proof.
$(i):$ By Theorem 2.8 and Remark 2.4,
$\Lambda_{1}(\Omega_{+})\geq\Lambda_{1}(\Omega_{+}^{*})=\left(\frac{|\Omega_{+}|}{|B_{1}|}\right)^{2/N}\Lambda_{1}(B_{1}).$
Similarly, $\Lambda_{1}(\Omega_{-})\geq\left(\frac{|\Omega_{-}|}{|B_{1}|}\right)^{2/N}\Lambda_{1}(B_{1}).$ Hence, by Lemma 2.13,
$\Lambda_{2}(\Omega)$ $>\max\{\Lambda_{1}(\Omega_{+}),\Lambda_{1}(\Omega_{-})\}\geq\max\left\{\left(\frac{|\Omega_{+}|}{|B_{1}|}\right)^{2/N}\Lambda_{1}(B_{1}),\left(\frac{|\Omega_{-}|}{|B_{1}|}\right)^{2/N}\Lambda_{1}(B_{1})\right\}$
$=\frac{\Lambda_{1}(B_{1})}{|B_{1}|^{2/N}}\left(\max\{|\Omega_{+}|,|\Omega_{-}|\}\right)^{2/N}\geq\left(\frac{|\Omega|/2}{|B_{1}|}\right)^{2/N}\Lambda_{1}(B_{1})=\Lambda_{1}(B_{r(|\Omega|/2)}),$
where we used that $2\max\{|\Omega_{+}|,|\Omega_{-}|\}\geq|\Omega_{+}|+|\Omega_{-}|=|\Omega|$.
$(ii):$ For simplicity, we take $c:=2|B_{1}|$. A scaling argument using Remark 2.4 gives the statement for any $c>0$. Let $\phi_{1}$ denote the eigenfunction associated to the first eigenvalue $\Lambda_{1}:=\Lambda_{1}(B_{1})$ with $\int_{\mathbb{R}^{N}}Q_{B_{1}}\phi_{1}^{2}=1$. It is clear that $\phi_{1}(\,\cdot\,-ne_{1})$ and $\phi_{1}(\,\cdot\,+ne_{1})$ are the eigenfunctions associated to the first eigenvalues $\Lambda_{1}(B_{1}(ne_{1}))$ and $\Lambda_{1}(B_{1}(-ne_{1}))$, respectively, and $\Lambda_{1}=\Lambda_{1}(B_{1}(ne_{1}))=\Lambda_{1}(B_{1}(-ne_{1})).$ Let $u_{n}(x):=\phi_{1}(x+ne_{1})-\phi_{1}(x-ne_{1})$ for $x\in\mathbb{R}^{N}$. Using that $\int_{\mathbb{R}^{N}}Q_{B_{1}}\phi_{1}^{2}=1$ and that $\phi_{1}$ is uniformly bounded in $\mathbb{R}^{N}$ and decays exponentially (by Theorem 2.5), it is not hard to verify that
$\int_{\mathbb{R}^{N}}Q_{\Omega_{n}}u_{n}^{2}=2+o(1)$ (2.16)
and
$\int_{\mathbb{R}^{N}}\nabla\phi_{1}(x+ne_{1})\cdot\nabla\phi_{1}(x-ne_{1})=\Lambda_{1}(B_{1})\int_{\mathbb{R}^{N}}Q_{B_{1}(-ne_{1})}\phi_{1}(x+ne_{1})\phi_{1}(x-ne_{1})=o(1)$ (2.17)
as $n\to\infty.$ Let $f_{n}$ be the eigenfunction associated to the first eigenvalue $\alpha_{n}:=\Lambda_{1}(\Omega_{n})$ with $\int_{\mathbb{R}^{N}}Q_{\Omega_{n}}f_{n}^{2}=1.$ Since the first eigenvalue is simple,
$\phi_{1}(x)=\phi_{1}(\sigma(x))\quad\text{and}\quad f_{n}(x)=f_{n}(\sigma(x))\quad\text{ for all }x\in\mathbb{R}^{N},\qquad\text{where \ }\sigma(x):=(-x_{1},...,x_{N}).$ (2.18)
Then, changing variables,
$\int_{\mathbb{R}^{N}}Q_{\Omega_{n}}(x)f_{n}(x)\phi_{1}(x-ne_{1})\,\mathrm{d}x=\int_{\mathbb{R}^{N}}Q_{\Omega_{n}}(x)f_{n}(x)\phi_{1}(\sigma(x)-ne_{1})\,\mathrm{d}x=\int_{\mathbb{R}^{N}}Q_{\Omega_{n}}(x)f_{n}(x)\phi_{1}(x+ne_{1})\,\mathrm{d}x$
and, as a consequence,
$\int_{\mathbb{R}^{N}}\nabla f_{n}\cdot\nabla u_{n}$ $=\alpha_{n}\left(\int_{\mathbb{R}^{N}}Q_{\Omega_{n}}(x)f_{n}(x)\phi_{1}(x-ne_{1})\,\mathrm{d}x-\int_{\mathbb{R}^{N}}Q_{\Omega_{n}}(x)f_{n}(x)\phi_{1}(x+ne_{1})\,\mathrm{d}x\right)=0.$
Hence, $f_{n}$ and $u_{n}$ are orthogonal in $D^{1,2}(\mathbb{R}^{N})$. If $\psi_{n}:=u_{n}/\sqrt{\int_{\mathbb{R}^{N}}Q_{\Omega_{n}}u_{n}^{2}},$ then, using statement $(i)$, (2.16) and (2.17),
$\Lambda_{1}(B_{1})<\Lambda_{2}(\Omega_{n})=\inf_{\begin{subarray}{c}v\in S_{\Omega_{n}}\\ v\perp f_{n}\end{subarray}}\|v\|^{2}\leq\|\psi_{n}\|^{2}=\frac{2\int_{\mathbb{R}^{N}}|\nabla\phi_{1}|^{2}}{2+o(1)}+o(1)=\Lambda_{1}(B_{1})+o(1)\qquad\text{ as }n\to\infty,$
as claimed. ∎
Summarizing, we obtain Theorem 1.6.
Proof of Theorem 1.6.
Part 1. This is an immediate consequence of Proposition 2.6 and Lemma 2.7, which establish the nodal properties and sign-changing behavior of the eigenfunctions.
Part 2. The sharp decay estimates are proven in Theorem 2.5.
Part 3. The Faber–Krahn-type inequality is established in Theorem 2.8.
Part 4. The Hong–Krahn–Szegö-type inequality is proved in Theorem 2.14.
Part 5. The symmetry properties of the eigenfunctions are derived in Propositions 2.9 and 2.12, which treat the radial symmetry of the first eigenfunction and the foliated Schwarz symmetry of the second one, respectively.
2.3 Nonexistence of negative eigenvalues
The following Pohožaev-type identity was proved in [5, Theorem 3.1].
Theorem 2.15.
Let $\Theta$ be a (possibly unbounded) smooth open subset of $\mathbb{R}^{N}$ and $f:\mathbb{R}\to\mathbb{R}$ be a continuous function. Set $F(s):=\int_{0}^{s}f$. If $u\in C^{1}(\mathbb{R}^{N})\cap C^{2}(\Theta)\cap C^{2}(\mathbb{R}^{N}\smallsetminus\overline{\Theta})\cap D^{1,2}(\mathbb{R}^{N})$ is a solution to
$\Delta u=Q_{\Theta}f(u)$ (2.19)
such that $F(u)\in L^{1}(\mathbb{R}^{N})$, then $u$ satisfies
$\frac{1}{2^{*}}\int_{\mathbb{R}^{N}}|\nabla u|^{2}-\int_{\mathbb{R}^{N}}Q_{\Theta}F(u)+\frac{2}{N}\int_{\partial\Theta}F(u)\,\zeta\cdot\nu_{\Theta}\,d\zeta=0,$
where $\nu_{\Theta}$ denotes the exterior unit normal to $\Theta$.
We apply the Pohožaev identity to prove our nonexistence result, Theorem 1.7, stated in the introduction.
Proof of Theorem 1.7.
Applying Theorem 2.15, we have that
$0$ $=\frac{1}{2^{*}}\int_{\mathbb{R}^{N}}|\nabla u|^{2}-\frac{\Lambda}{2}\int_{\mathbb{R}^{N}}Q_{\Omega}u^{2}+\frac{\Lambda}{N}\int_{\partial\Omega}u^{2}\zeta\cdot\nu_{\Omega}\,d\zeta=\left(\frac{1}{2^{*}}-\frac{1}{2}\right)\int_{\mathbb{R}^{N}}|\nabla u|^{2}+\frac{\Lambda}{N}\int_{\partial\Omega}u^{2}\,\zeta\cdot\nu_{\Omega}\,d\zeta.$
This implies that $\frac{\Lambda}{N}\int_{\partial\Omega}u^{2}\,\zeta\cdot\nu_{\Omega}\,d\zeta>0$, and the claim follows. ∎
3 The almost linear problem
In this section we consider the nonlinear problem, labelled (1.1) in the introduction,
$\begin{cases}-\Delta u=Q_{\Omega}|u|^{p-2}u,\\ u\in X^{p}:=D^{1,2}(\mathbb{R}^{N})\cap L^{p}(\mathbb{R}^{N}).\end{cases}$
We study the behavior of least energy solutions for $p$ close to $2$.
3.1 Asymptotic behavior of ground states
We adapt the approach followed in [21]. For $p\in(1,2^{*})$ consider the Banach space $X^{p}:=D^{1,2}(\mathbb{R}^{N})\cap L^{p}(\mathbb{R}^{N})$ equipped with the norm
$\|u\|_{X^{p}}:=(\|u\|^{2}+|u|_{p}^{2})^{\frac{1}{2}},$
where $|\cdot|_{p}$ denotes the norm in $L^{p}(\mathbb{R}^{N})$. Set
$\mathcal{S}_{\Omega,p}:=\left\{v\in X^{p}:\int_{\mathbb{R}^{N}}Q_{\Omega}(x)|v|^{p}=1\right\},$
and consider the minimization problem
$\alpha_{p}:=\inf_{v\in\mathcal{S}_{\Omega,p}}\|v\|^{2}.$ (3.1)
Lemma 3.1.
$0<a_{0}\leq\alpha_{p}\leq a_{1}<\infty$ for all $p\in(1,2^{*})$.
Proof.
Let $u\in\mathcal{S}_{\Omega,p}$. Since $\Omega$ is bounded, using Hölder’s and Sobolev’s inequalities we obtain
$1$ $=\int_{\mathbb{R}^{N}}Q_{\Omega}(x)|u|^{p}\leq\int_{\Omega}|u|^{p}\leq|\Omega|^{\frac{2^{*}-p}{2^{*}}}\left(\int_{\Omega}|u|^{2^{*}}\right)^{p/2^{*}}$
$\leq|\Omega|^{\frac{2^{*}-p}{2^{*}}}\left(\int_{\mathbb{R}^{N}}|u|^{2^{*}}\right)^{p/2^{*}}\leq|\Omega|^{\frac{2^{*}-p}{2^{*}}}C\left(\int_{\mathbb{R}^{N}}|\nabla u|^{2}\right)^{p/2},$
where $C=C(N)>0$. Therefore, $a_{0}:=\min_{p\in[1,2^{*}]}(|\Omega|^{\frac{(p-2^{*})2}{2^{*}p}}C^{-2/p})\leq\|u\|^{2}$ for every $u\in\mathcal{S}_{\Omega,p}$ and $p\in(1,2^{*})$. This shows that $\alpha_{p}\geq a_{0}>0$.
To prove the other inequality, fix $\rho\in C_{c}^{\infty}(\Omega)$ such that $\rho\neq 0$. Then, the function $f:[1,\infty)\to D^{1,2}(\mathbb{R}^{N})$, given by $f(p):=\frac{\rho}{\left(\int_{\mathbb{R}^{N}}Q_{\Omega}|\rho|^{p}\right)^{1/p}}$, is continuous and $f(p)\in\mathcal{S}_{\Omega,p}$. Hence
$\alpha_{p}\leq\max_{p\in[1,2^{*}]}\|f(p)\|^{2}=:a_{1}<\infty\qquad\text{for all}\;p\in[1,2^{*}],$ (3.2)
as claimed. ∎
Theorem 3.2.
[…]
Proof.
The proof is similar to that of Lemma  2.1. It is given in [14, Lemma 2.1]. ∎
A solution $u_{p}:=\alpha_{p}^{\frac{1}{p-2}}v_{p}$ to (1.1) such that $v_{p}$ is a minimizer for $\alpha_{p}$ is called a least energy solution. These solutions are the minimizers on the space $X^{p}$ of the energy functional given by
$J(u):=\frac{1}{2}\int_{\mathbb{R}^{N}}|\nabla u|^{2}-\frac{1}{p}\int_{\mathbb{R}^{N}}Q_{\Omega}|u|^{p}$
if $p\in(1,2)$, and they are the minimizers of $J$ on the Nehari manifold $\mathcal{N}$, defined in [6, (3.2)], if $p\in(2,2^{*})$.
Lemma 3.3.
Every solution to (1.1) belongs to $W_{loc}^{2,s}(\mathbb{R}^{N})\cap C_{loc}^{1,\alpha}(\mathbb{R}^{N})\cap C_{loc}^{\infty}(\Omega)\cap C_{loc}^{\infty}(\mathbb{R}^{N}\setminus\overline{\Omega})$ for all $s\in[1,\infty)$ and $\alpha\in(0,1)$.
Proof.
See [6, Lemma 2.4]. ∎
Theorem 3.4.
For $p\in(1,2^{*})$, let $v_{p}\in\mathcal{S}_{\Omega,p}$ be a nonnegative minimizer of (3.1). Then $v_{p}\to\phi_{1}$ in $D^{1,2}(\mathbb{R}^{N})$ as $p\to 2$, where $\phi_{1}$ is the positive eigenfunction of (1.2) with eigenvalue $\Lambda_{1}$ such that $\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{2}=1$. Moreover,
$\lim_{p\to 2}\left(\frac{\Lambda_{1}}{\alpha_{p}}\right)^{\frac{1}{2-p}}=e^{-\frac{1}{2}\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{2}\ln(\phi_{1}^{2})}.$ (3.3)
Proof.
Fix $\delta>0$ so that $2-\delta,2+\delta]\subset(1,2^{*})$. Let $p_{n}\in(1,2^{*})$ be such that $p_{n}\to 2$ and $v_{n}\in S_{\Omega,p_{n}}$ be such that $\alpha_{p_{n}}=\|v_{n}\|^{2}$. By Lemma 3.1 we have that $(v_{n})$ is uniformly bounded in $D^{1,2}(\mathbb{R}^{N})$. Hence, passing to a subsequence, $v_{n}\rightharpoonup v$ weakly in $D^{1,2}(\mathbb{R}^{N})$, $v_{n}\to v$ in $L_{loc}^{q}(\mathbb{R}^{N})$ for every $q\in[2-\delta,2+\delta]$, and $v_{n}\to v$ a.e. in $\mathbb{R}^{N}$. Then, $v\geq 0$. We claim that
$\lim_{n\to\infty}\int_{\Omega}v_{n}^{p_{n}}=\lim_{n\to\infty}\int_{\Omega}v_{n}^{2}=\int_{\Omega}v^{2}<\infty.$ (3.4)
For $n$ large enough, we have that $2s+(1-s)p_{n}\in(2-\delta/2,2+\delta/2)$ for all $s\in(0,1)$. Since $\lim_{x\to 0^{+}}x^{\delta/2}\ln(x)=0$ and $\lim_{x\to\infty}\ln(x)/x^{\delta}=0$, there exists $C=C(\delta)>0$ such that $x^{\delta/2}|\ln(x)|\leq Cx^{\delta}$ for all $x\in(0,\infty)$, and
$\left|\int_{\Omega}v_{n}^{p_{n}}-\int_{\Omega}v_{n}^{2}\right|$ $\leq|p_{n}-2|\int_{\Omega}\int_{0}^{1}v_{n}^{2s+(1-s)p_{n}}|\ln(v_{n})|\,\mathrm{d}s\,\mathrm{d}x$
$\leq|p_{n}-2|\int_{\Omega}(v_{n}^{2-\delta/2}+v_{n}^{2+\delta/2})|\ln(v_{n})|\,\mathrm{d}x\leq C|p_{n}-2|\int_{\Omega}(v_{n}^{2}+v_{n}^{2+\delta})\,\mathrm{d}x=o(1)$
as $n\to\infty$, which yields (3.4). Furthermore, as $v_{n}\in\mathcal{S}_{\Omega,p_{n}}$, Fatou’s lemma and (3.4) yield
$\int_{\mathbb{R}^{N}\smallsetminus\Omega}v^{2}\leq\liminf_{n\to\infty}\int_{\mathbb{R}^{N}\smallsetminus\Omega}|v_{n}|^{p_{n}}=\lim_{n\to\infty}\int_{\Omega}|v_{n}|^{p_{n}}-1=\int_{\Omega}v^{2}-1.$
This shows that $v\in L^{2}(\mathbb{R}^{N})$ and that $1\leq\int_{\mathbb{R}^{N}}Q_{\Omega}v^{2}$. In particular, $v\neq 0$. Hence,
$\liminf_{n\to\infty}\alpha_{p_{n}}=\liminf_{n\to\infty}\|v_{n}\|^{2}\geq\|v\|^{2}\geq\frac{\|v\|^{2}}{\int_{\mathbb{R}^{N}}Q_{\Omega}v^{2}}\geq\Lambda_{1}.$ (3.5)
On the other hand, by Theorem 2.5 and Lemma 3.3, the first positive eigenfunction $\phi_{1}$ of (1.2) with $\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{2}=1$, satisfies $\phi_{1}\in L^{p}(\mathbb{R}^{N})$ for every $p\in[1,\infty)$. Fixing $\alpha\in(0,\sqrt{\Lambda_{1}})$, there exists $C>0$ such that $\phi_{1}^{p_{n}}(x)\leq Ce^{-\alpha|x|}$ for large enough $n$, and the dominated convergence theorem implies that $\lim_{p\to 2}\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{p}=\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{2}=1$. Therefore,
$\Lambda_{1}=\|\phi_{1}\|^{2}=\lim_{n\to\infty}\frac{\|\phi_{1}\|^{2}}{\Big(\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{p_{n}}\Big)^{2/p_{n}}}\geq\limsup_{n\to\infty}\alpha_{p_{n}}.$ (3.6)
From (3.5) and (3.6) we get $\int_{\mathbb{R}^{N}}Q_{\Omega}v^{2}=1$, $\|v\|^{2}=\Lambda_{1}$, $v_{n}\to v$ strongly in $D^{1,2}(\mathbb{R}^{N})$, and
$\lim_{n\to\infty}\alpha_{p_{n}}=\Lambda_{1}.$ (3.7)
Since $\text{dim}(E_{1})=1$, we have that $v=\phi_{1}$.
It remains to show (3.3). Since $u_{n}:=\alpha_{p_{n}}^{\frac{1}{p_{n}-2}}v_{n}$ solves (1.1), we have that
$\int_{\mathbb{R}^{N}}\Lambda_{1}Q_{\Omega}\phi_{1}v_{n}=\int_{\mathbb{R}^{N}}\nabla\phi_{1}\cdot\nabla v_{n}=\int_{\mathbb{R}^{N}}\alpha_{p_{n}}Q_{\Omega}v_{n}^{p-1}\phi_{1}.$
Hence,
$0$ $=\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}v_{n}\left(\frac{\alpha_{p_{n}}}{\Lambda_{1}}v_{n}^{p_{n}-2}-1\right)=\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}v_{n}\int_{0}^{1}\frac{d}{ds}\Big(\Big(\frac{\alpha_{p_{n}}}{\Lambda_{1}}\Big)^{\frac{1}{p_{n}-2}}v_{n}\Big)^{s(p_{n}-2)}\,\mathrm{d}s\,\mathrm{d}x$
$=(p_{n}-2)\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}v_{n}\int_{0}^{1}\ln\Big(\Big(\frac{\alpha_{p_{n}}}{\Lambda_{1}}\Big)^{\frac{1}{p_{n}-2}}v_{n}\Big)\Big(\Big(\frac{\alpha_{p_{n}}}{\Lambda_{1}}\Big)^{\frac{1}{p_{n}-2}}v_{n}\Big)^{s(p_{n}-2)}\,\mathrm{d}s\,\mathrm{d}x$
$=(p_{n}-2)\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}v_{n}\left(\frac{1}{p_{n}-2}\ln\Big(\frac{\alpha_{p_{n}}}{\Lambda_{1}}\Big)+\ln(v_{n})\right)\int_{0}^{1}\left(\frac{\alpha_{p_{n}}}{\Lambda_{1}}\right)^{s}v_{n}^{s(p_{n}-2)}\,\mathrm{d}s\,\mathrm{d}x.$
Since $v_{n}\to\phi_{1}$ in $D^{1,2}(\mathbb{R}^{N})$ there exists $h\in L^{2^{*}}(\mathbb{R}^{N})$ such that, after passing to a subsequence, $v_{n}\leq h$ a.e. in $\mathbb{R}^{N}$ (see, for instance, [27, Lemma A.1]). By (3.7), there exists $M>1$ such that
$\left(\frac{\alpha_{p_{n}}}{\Lambda_{1}}\right)^{s}\leq M\quad\text{for all}\;s\in(0,1)\;\text{and}\;n\in\mathbb{N}.$
Since $p_{n}\to 2,$ we have that $s(p_{n}-2)+1\in(1/2,2)$ for $n$ large enough. By Theorem 2.5, there exists $C>0$ such that $\phi_{1}\leq Ce^{-\sqrt{\Lambda_{1}}|x|}$ for all $x\in\mathbb{R}^{N}$. It follows that
$0<v_{n}\phi_{1}\left(\frac{\alpha_{p_{n}}}{\Lambda_{1}}\right)^{s}v_{n}^{s(p_{n}-2)}\leq M\phi_{1}h^{s(p_{n}-2)+1}\leq MCe^{-\sqrt{\Lambda_{1}}|x|}(h^{1/2}+h^{2}).$
On the other hand, fixing $\varepsilon\in(0,\frac{1}{2})$ such that $2+\varepsilon<2^{*}$, we have that
$\Big|v_{n}\phi_{1}\ln(v_{n})\Big(\frac{\alpha_{p_{n}}}{\Lambda_{1}}\Big)^{s}v_{n}^{s(p_{n}-2)}\Big|$ $\leq MC_{\varepsilon}\phi_{1}v_{n}^{s(p_{n}-2)+1+\varepsilon}\leq MC_{\varepsilon}\phi_{1}(h^{\frac{1}{2}+\varepsilon}+h^{2+\varepsilon})$
$\leq MC_{\varepsilon}Ce^{-\sqrt{\Lambda_{1}}|x|}(h^{\frac{1}{2}+\varepsilon}+h^{2+\varepsilon}),$
where $C_{\varepsilon}>0$ is such that $|x\ln(|x|)|\leq C_{\varepsilon}|x|^{1+\varepsilon}$ for all $x\in\mathbb{R}$. Then, by dominated convergence theorem,
$\int_{\mathbb{R}^{N}}\int_{0}^{1}Q_{\Omega}v_{n}\phi_{1}\Big(\frac{\alpha_{p_{n}}}{\Lambda_{1}}\Big)^{s}v_{n}^{s(p_{n}-2)}\,\mathrm{d}s\,\mathrm{d}x>0$
for $n$ large enough. Moreover,
$\frac{\ln\big(\frac{\alpha_{p_{n}}}{\Lambda_{1}}\big)}{p_{n}-2}=-\frac{\int_{\mathbb{R}^{N}}\int_{0}^{1}Q_{\Omega}v_{n}\phi_{1}\ln(v_{n})\left(\frac{\alpha_{p_{n}}}{\Lambda_{1}}\right)^{s}v_{n}^{s(p_{n}-2)}\,\mathrm{d}s\,\mathrm{d}x}{\int_{\mathbb{R}^{N}}\int_{0}^{1}Q_{\Omega}v_{n}\phi_{1}\left(\frac{\alpha_{p_{n}}}{\Lambda_{1}}\right)^{s}v_{n}^{s(p_{n}-2)}\,\mathrm{d}s\,\mathrm{d}x}$
and, using that $\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{2}=1$,
$\ln\Big(\Big(\frac{\alpha_{p_{n}}}{\Lambda_{1}}\Big)^{\frac{1}{p_{n}-2}}\Big)=\frac{\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{2}\ln(\phi_{1})}{\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{2}}+o(1)=-\frac{1}{2}\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{2}\ln(\phi_{1}^{2})+o(1),$
as claimed. ∎
Theorem 1.4 stated in the introduction is now an immediate consequence of the previous result.
3.2 Uniqueness of minimizers and nondegeneracy of positive solutions
In this section we prove Theorem 1.1. The proof follows the strategy used in [13], but several adaptations are needed in order to control the lack of compactness at spatial infinity.
Consider the linearized problem at a positive solution $u$ to problem (1.1) given by
$\Delta h=Q_{\Omega}(p-1)u^{p-2}h\qquad\text{ in }\mathbb{R}^{N}.$ (3.8)
A solution $u$ to (1.1) is said to be nondegenerate if (3.8) has only the trivial solution $h=0$, i.e. if $\mu=0$ is not an eigenvalue of the linearized operator $L_{u}:=-\Delta-Q_{\Omega}(p-1)u^{p-2}$. The spectrum and the associated eigenfunctions for $L_{u}$ can be obtained analogously to Section 2.
We first need to prove some auxiliary results. We begin with a nonexistence result.
Lemma 3.5.
Let $U\subset\mathbb{R}^{N}$ be an open bounded set such that $\Lambda_{1}(U)=1.$ If $\Omega\subset\mathbb{R}^{N}$ is and open set such that $U\subset\Omega$, $U\neq\Omega,$ then the problem
$-\Delta u=Q_{\Omega}u\quad\text{ in }\mathbb{R}^{N},\qquad u>0\quad\text{ in }\mathbb{R}^{N},$ (3.9)
does not admit a bounded positive solution. As a consequence, if $\Omega=\mathbb{R}^{N}_{+}:=\{x\in\mathbb{R}^{N}\>:\>x_{N}>0\}$ is a halfspace or if $\Omega=\mathbb{R}^{N}$, then (3.9) does not admit a bounded positive solution.
Proof.
Assume, by contradiction, that $u$ is a bounded positive solution to problem (3.9). By maximum principles, arguing as in Theorem 2.5, we have that $u$ must have exponential decay and, therefore, $u\in X^{p}.$ Let $\phi_{U}$ denote the eigenfunction associated to the first eigenvalue $\Lambda_{1}(U)=1$. Testing (3.9) with $\phi_{U}$,
$\int_{\mathbb{R}^{N}}Q_{U}u\phi_{U}=\int_{\mathbb{R}^{N}}u(-\Delta\phi_{U})=\int_{\mathbb{R}^{N}}(-\Delta u)\phi_{U}=\int_{\mathbb{R}^{N}}Q_{\Omega}u\phi_{U},$
and, therefore,
$\int_{\mathbb{R}^{N}}(Q_{\Omega}-Q_{U})u\phi_{U}=0.$ (3.10)
On the other hand, since $U\subset\Omega,$ we have $Q_{\Omega}-Q_{U}\geq 0$ in $\mathbb{R}^{N}$ and $Q_{\Omega}-Q_{U}=2>0$ in $\U\Omega\backslash$ U, which contradicts (3.10), because $u\phi_{U}>0$ in $\mathbb{R}^{N}$.
In particular, if $\Omega=\mathbb{R}^{N}_{+}$ or $\Omega=\mathbb{R}^{N}$, then (3.9) does not admit a positive solution, because $\mathbb{R}^{N}_{+}$ contains the ball $B_{r}(re_{N})$, where $r>0$ is such that $\Lambda_{1}(B_{r}(re_{N}))=1.$
Lemma 3.6.
Let $(p_{n})$ be a sequence in $(2,2^{*})$ such that $p_{n}\to 2$, $u_{n}$ be a positive solution to (1.1) with $p=p_{n}$, and let $M_{n}:=|u_{n}|_{\infty}$. Then,
$M_{n}^{p_{n}-2}\to\Lambda_{1}(\Omega)\qquad\text{and}\qquad\frac{u_{n}}{M_{n}}\to\phi_{1}\text{ \ in \ }C^{2,\alpha}_{loc}(\Omega)\cap C^{2,\alpha}_{loc}(\mathbb{R}^{N}\backslash\Omega)\cap C^{1,\alpha}_{loc}(\mathbb{R}^{N})\qquad\text{as \ }n\to\infty$
for some $\alpha\in(0,1)$, where $\Lambda_{1}(\Omega)$ and $\phi_{1}$ denote, respectively, the first eigenvalue and eigenfunction normalized so that $|\phi_{1}|_{\infty}=1$.
Proof.
Step 1. We show that $M_{n}^{p_{n}-2}$ is bounded.
By contradiction, assume that $M_{n}^{p_{n}-2}\to\infty$ and let $x_{n}\in\overline{\Omega}$ be such that $M_{n}=u_{n}(x_{n})$ (the fact that $x_{n}\in\overline{\Omega}$ follows easily from (1.1), since $\Delta u_{n}(x_{n})\geq 0$ whenever $x_{n}\not\in\partial\Omega,$ because $x_{n}$ is a global maximum). Define
$w_{n}(y):=\frac{1}{M_{n}}u_{n}\left(\mu_{n}y+x_{n}\right)$
where $\mu_{n}:=M_{n}^{\frac{2-p_{n}}{2}}\to 0$ as $n\to\infty$. Then $0\leq w_{n}\leq 1$, $w_{n}(0)=1$, and
$-\Delta w_{n}=Q_{\Omega_{n}}w_{n}^{p_{n}-1}=:f_{n}\quad$ $\text{ in }\mathbb{R}^{N},$ (3.11)
where $\Omega_{n}:=\left\{y\in\mathbb{R}^{N}:\,\mu_{n}y+x_{n}\in\Omega\right\}$. Up to subsequences,
$\text{ either }\quad\operatorname{dist}\left(x_{n},\,\partial\Omega\right)\mu_{n}^{-1}\rightarrow+\infty\quad\text{ or }\quad\operatorname{dist}\left(x_{n},\,\partial\Omega\right)\mu_{n}^{-1}\rightarrow\rho\geq 0.$ (3.12)
Assume the first case holds, so that $\Omega_{n}\rightarrow\mathbb{R}^{N}$ as $n\rightarrow+\infty$. For every $R>0$ we claim that there exists $n_{R}\in\mathbb{N}$ and $C>0$ such that
$w_{n}\in C^{2,\alpha}(\overline{B}_{R})\ \mbox{ and }\ \|w_{n}\|_{C^{2,\alpha}(\overline{B}_{R})}\leq C\quad\text{ for all }n\geq n_{R}$ (3.13)
and all $\alpha\in(0,1)$. In order to prove (3.13) let us fix $R>0$. Then, since $\Omega_{n}\rightarrow\mathbb{R}^{N}$ as $n\rightarrow+\infty$, there exists $n_{R}\in\mathbb{N}$ such that $\overline{B}_{4R}\subset\Omega_{n}$ for any $n\geq n_{R}$. Since $0\leq w_{n}\leq 1$ in $\mathbb{R}^{N}$, then $|f_{n}|\leq 1$ in $\mathbb{R}^{N}$. In particular, by elliptic regularity, $w_{n}\in C^{1,\alpha}(\overline{B}_{2R})$ for any $\alpha\in(0,1)$ and $\|w_{n}\|_{C^{1,\alpha}(B_{2R})}\leq C$ for some constant $C=C(N,\alpha,R)>0$. Then $f_{n}\in C^{\alpha}(B_{2R})$ with uniform $C^{\alpha}$-bound and so (3.13) follows by elliptic regularity.
By (3.13), using the Arzelà-Ascoli theorem and a diagonal argument we obtain that there exists a function $w\in C_{loc}^{2,\frac{\alpha}{2}}(\mathbb{R}^{N})$ such that, passing to a subsequence, $w_{n}\rightarrow w$ in $C^{2,\frac{\alpha}{2}}_{loc}(\mathbb{R}^{N})$. Passing (3.11) to the limit, we see that $w$ solves $\Delta w=w$ in $\mathbb{R}^{N}$ pointwisely. Furthermore, $0\leq w\leq 1$, so $w$ is bounded in $\mathbb{R}^{N}$ and $w>0$ in $\mathbb{R}^{N}$ by the maximum principle. However, by Lemma 3.5, no such $w$ can exist and we have reached a contradiction.
If the second case in (3.12) holds then we may assume $x_{n}\rightarrow x_{0}\in\partial\Omega.$ With no loss of generality assume that $\nu(x_{0})=-e_{N}$, where $\nu$ denotes the exterior unit normal vector at $\partial\Omega$. Let
$w_{n}(y):=\frac{u_{n}\left(\mu_{n}y+\xi_{n}\right)}{M_{n}},$
where $\xi_{n}\in\partial\Omega$ is the closest point to $x_{n}$ on $\partial\Omega$. Let $D_{n}:=\left\{y\in\mathbb{R}^{N}:\mu_{n}y+\xi_{n}\in\Omega\right\}$ and observe that
$0\in\partial D_{n}$ (3.14)
and $D_{n}\rightarrow\mathbb{R}^{N}_{+}:=\left\{y\in\mathbb{R}^{N}:y_{N}>0\right\}$ as $n\rightarrow+\infty$. It also follows that $w_{n}$ satisfies that
$-\Delta w_{n}=Q_{D_{n}}w_{n}^{p_{n}-1}\quad\text{ in }\mathbb{R}^{N}.$ (3.15)
Moreover, setting
$y_{n}:=\frac{x_{n}-\xi_{n}}{\mu_{n}}\ (\in\overline{D_{n}})$
we have that $\left|y_{n}\right|=\operatorname{dist}\left(x_{n},\,\partial\Omega\right)\mu_{n}^{-1}$ and $w_{n}(y_{n})=1$. Now, arguing similarly as in the first case, we obtain that $w_{n}\rightarrow w$ in $C^{2,\frac{\alpha}{2}}_{loc}(\mathbb{R}^{N}_{+})\cup C^{2,\frac{\alpha}{2}}_{loc}(-\mathbb{R}^{N}_{+})\cup C^{1,\frac{\alpha}{2}}_{loc}(\mathbb{R}^{N})$, where $w$ satisfies that $0\leq w\leq 1$ in $\mathbb{R}^{N}$, $w(y_{0})=1$, and $w$ is a bounded solution to
$-\Delta w=Q_{\mathbb{R}^{N}_{+}}w\qquad\text{ in }\mathbb{R}^{N}.$
By the maximum principle, $w>0$ in $\mathbb{R}^{N}$, which contradicts Lemma 3.5. This shows that $M_{n}^{p_{n}-2}$ is bounded and concludes the proof of Step 1.
Step 2. Conclusion.
$M_{n}^{p_{n}-2}$
is bounded by Step 1. Thus, up to a subsequence, $M_{n}^{p_{n}-2}\to\mu$. Let $z_{n}:=\frac{u_{n}}{M_{n}}$, then
$-\Delta z_{n}=Q_{\Omega}M_{n}^{p_{n}-2}\,z_{n}^{p_{n}-1}=:g_{n}\quad\text{ in }\mathbb{R}^{N},\qquad 0<z_{n}\leq 1\quad\text{ in }\mathbb{R}^{N}.$
Arguing as before, by elliptic regularity, $z_{n}$ converges to $z$ in $C_{loc}^{2,\frac{\beta}{2}}(\Omega)\cup C_{loc}^{2,\frac{\beta}{2}}(\mathbb{R}^{N}\backslash\Omega)\cup C^{1,\frac{\beta}{2}}(\mathbb{R}^{N})$; furthermore, $0\leq z\leq 1$ and it satisfies that
$-\Delta z=Q_{\Omega}\mu z\quad\text{ in }\mathbb{R}^{N},\qquad z\geq 0\quad\text{ in }\mathbb{R}^{N},$ (3.16)
Let $x_{n}\in\overline{\Omega}$ be such that $M_{n}=u_{n}(x_{n})$, then $1=\lim_{n\to\infty}z_{n}(x_{n})=z(\overline{x})$, for some $\overline{x}\in\overline{\Omega}$. By the maximum principle $z>0$ in $\mathbb{R}^{N}$ and then, using that only the first eigenfunction is positive and its eigenvalue is simple, we have that $\mu=\Lambda_{1}(\Omega)$, $z=\phi_{1}$. Note that any bounded solution to (3.16) must have the decay stated in Theorem 2.5 (by using a comparison argument as in the proof of this lemma) and, therefore, belongs to $H^{1}(\mathbb{R}^{N}).$ This ends the proof. ∎
Next, we show a uniform estimate for normalized positive solutions.
Lemma 3.7.
Fix $q\in(2,2\frac{N-1}{N-2})$. Let $(p_{n})$ be a sequence in $(2,q)$ such that $p_{n}\to 2$, $u_{n}$ be a positive solution to (1.1) with $p=p_{n}$, $M_{n}:=|u_{n}|_{\infty}$ and $z_{n}:=\frac{u_{n}}{M_{n}}$. There are $R>1$ and $C>0$ such that
$z_{n}(x)\leq C|x|^{-\frac{2}{q-2}}\qquad\text{if }|x|\geq R\text{ and for all }n\in\mathbb{N}.$ (3.17)
Proof.
[…]
$-\Delta\tau=-\tau^{q-1}\quad\text{in }\mathbb{R}^{N}\setminus\{0\}.$
Let $R>1$ be such that $\Omega\subset B_{R}(0)$ and define $\psi_{n}:=\tau-z_{n}$. Then, using that $u_{n}$ is a positive solution of (1.1) with $p=p_{n}$, we have that
$-\Delta\psi_{n}=-\tau^{q-1}+M_{n}^{p_{n}-2}z_{n}^{p_{n}-1}\quad\text{in }\mathbb{R}^{N}\setminus B_{R}(0).$
Let
$a_{n}(x):=\begin{cases}\frac{\tau(x)^{p_{n}-1}-z_{n}^{p_{n}-1}(x)}{\tau(x)-z_{n}(x)}M_{n}^{p_{n}-2},&\text{if }\tau(x)\neq z_{n}(x),\\ 0,&\text{if }\tau(x)=z_{n}(x).\end{cases}$
Observe that $a_{n}\geq 0$ and
$-\Delta\psi_{n}+a_{n}\psi_{n}=-\tau^{q-1}+M_{n}^{p_{n}-2}\tau^{p_{n}-1}\quad\text{in }\mathbb{R}^{N}\setminus B_{R}(0).$ (3.18)
Since $2<p_{n}<q$ and $M_{n}^{p_{n}-2}\to\Lambda_{1}(\Omega)$, note that
$-\tau^{q-1}+M_{n}^{p_{n}-2}\tau^{p_{n}-1}\geq\left(\frac{\Lambda_{1}}{2}-\tau^{q-p_{n}}\right)\tau^{p_{n}-1}\geq 0\quad\text{in }\mathbb{R}^{N}\setminus B_{R}(0)$ (3.19)
for $n$ and $R$ sufficiently large. Hence, by (3.18) and (3.19), there are $m\in\mathbb{N}$ and $R_{0}>R$ such that
$-\Delta\psi_{n}+a_{n}\psi_{n}\geq 0\qquad\text{in }\mathbb{R}^{N}\setminus B_{R_{0}}(0)\text{ \ for }n\geq m.$
On the other hand, by Theorem 2.5, $\phi_{1}(x)\leq C|x|^{-\frac{N-1}{2}}e^{-\sqrt{\Lambda_{1}(\Omega)}|x|}$. Hence, there is $R_{1}>R_{0}$ such that $\tau(x)-\phi_{1}(x)>0$ for all $x\in\mathbb{R}^{N}\setminus\overline{B_{R_{1}}(0)}$. Since, by Lemma 3.6, $z_{n}\to\phi$ in $C_{loc}^{1,\alpha}(\mathbb{R}^{N})$ for some $\alpha\in(0,1)$, there is $\tilde{m}\in\mathbb{N}$ such that
$\psi_{n}>0\qquad\text{ on $\partial B_{R_{1}}(0)$ \ for all $n\geq\tilde{m}$. }$ (3.20)
Since $z_{n}(x)\to 0$ as $|x|\to\infty$ (see Proposition 1.5) and (3.19) and (3.20) hold, it follows from the maximum principle that
$z_{n}(x)<\tau(x)\quad\text{for all }|x|\geq R_{1}$
and $n\geq\widetilde{m}$, as claimed. ∎
We are ready to prove our uniqueness and nondegeneracy result.
Proof of Theorem 1.1.
Let $(p_{n})$ be a subsequence of $(2,2^{*})$ such that $p_{n}\to 2$, $u_{n}$ be a solution to (1.1) with $p=p_{n}$, and $M_{n}:=|u_{n}|_{\infty}$.
Step 1. Nondegeneracy of positive solutions: Assume, by contradiction, that there exists a non-trivial solution $h_{n}$ to the linearized problem (3.8) with $p=p_{n}>2$ and $p_{n}\to 2$, namely,
$-\Delta h_{n}=Q_{\Omega}(p_{n}-1)u_{n}^{p_{n}-2}h_{n}\ =:g_{n}\quad{\text{ in }}\mathbb{R}^{N}.$ (3.21)
We may assume that $|h_{n}|_{\infty}=1$. Using Lemma 3.6,
$|g_{n}|_{\infty}\leq 2|u_{n}|^{p_{n}-2}_{\infty}=2M_{n}^{p_{n}-2}\leq C.$
So, by elliptic regularity, $h_{n}\to h$ in $C^{1,\alpha}_{loc}(\mathbb{R}^{N})$. Furthermore, again by elliptic regularity, $h_{n}\in C^{\infty}(\Omega)\cap C^{1}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$ (the smoothness follows by a standard bootstrap procedure). Taking $h_{n}$ as a test function, by Lemma 3.6, we derive that
$\|h_{n}\|^{2}$ $=\int_{\mathbb{R}^{N}}Q_{\Omega}(p_{n}-1)u_{n}^{p_{n}-2}h_{n}^{2}\leq(p_{n}-1)\int_{\Omega}u_{n}^{p_{n}-2}h_{n}^{2}\leq 2M_{n}^{p_{n}-2}|\Omega||h_{n}|_{\infty}^{2}=2M_{n}^{p_{n}-2}|\Omega|\leq C.$
Hence, up to a subsequence, $h_{n}$ converges to $h$ weakly in $D^{1,2}(\mathbb{R}^{N})$ and strongly in $L_{loc}^{2}(\mathbb{R}^{N})$. Letting $n\to\infty$ in the weak formulation of (3.21) we obtain that $h$ is a weak solution of
$-\Delta h=Q_{\Omega}\Lambda_{1}(\Omega)h\quad{\text{ in }}\mathbb{R}^{N},$
because, by Lemma 3.6, one has that
$(p_{n}-1)u_{n}^{p_{n}-2}=(p_{n}-1)M_{n}^{p_{n}-2}\left(\frac{u_{n}}{M_{n}}\right)^{p_{n}-2}=\Lambda_{1}(\Omega)+o(1)\quad\text{pointwisely in $\mathbb{R}^{N}$ as }n\to\infty$ (3.22)
and
$|(p_{n}-1)u_{n}^{p_{n}-2}|_{\infty}\leq\Lambda_{1}(\Omega)+1\qquad\text{ for all }n\in\mathbb{N}.$ (3.23)
Note that both $h_{n}$ and $\phi_{1}$ achieve their supremum in $\overline{\Omega}$. Indeed, by (uniform) local regularity estimates and the fact that $h_{n}\in D^{1,2}(\mathbb{R}^{N})$, we have that $h_{n}$ decays uniformly to 0 at infinity. Let $(x_{n})$ be the points in $\mathbb{R}^{N}$ where $h_{n}(x_{n})=|h_{n}|_{\infty}=1$, and assume, by contradiction, that $x_{n}\not\in\overline{\Omega}$. Then, $0\leq-\Delta h_{n}(x_{n})=-(p_{n}-1)u_{n}(x_{n})^{p_{n}-1}h_{n}(x_{n})<0$, which yields a contradiction. That $\phi_{1}$ achieves its supremum in $\overline{\Omega}$ can be shown similarly.
Then, by the local convergence, $h=\phi_{1}$, where $\phi_{1}$ is the eigenfunction associated to the first eigenvalue $\Lambda_{1}(\Omega)$ such that $|\phi_{1}|_{\infty}=1$. Note that, integrating by parts,
$0=\int_{\mathbb{R}^{N}}h_{n}(-\Delta u_{n})-u_{n}(-\Delta h_{n})\,dx=\int_{\mathbb{R}^{N}}Q_{\Omega}u_{n}^{p_{n}-1}h_{n}\,dx=(2-p_{n})M_{n}^{p_{n}-1}\int_{\mathbb{R}^{N}}Q_{\Omega}z_{n}^{p_{n}-1}h_{n}\,dx.$ (3.24)
Using Lemmas 3.63.7, and the fact that $|h_{n}|_{\infty}=1$, we can use dominated convergence to obtain that
$0=\lim_{n\to\infty}\int_{\mathbb{R}^{N}}Q_{\Omega}z_{n}^{p_{n}-1}h_{n}\,dx=\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{2}\,dx>0.$
This yields a contradiction and the claim follows.
Step 2. Uniqueness of minimizers.
Arguing by contradiction, assume that there is a sequence $(p_{n})$ in $(2,2^{*})$ such that $p_{n}\to 2$ and, for each $n$, there are two different least energy solutions $u_{n}=\alpha_{n}^{\frac{1}{p_{n}-2}}v_{n}$ and $\hat{u}_{n}=\alpha_{n}^{\frac{1}{p_{n}-2}}\hat{v}_{n}$, where $\alpha_{n}:=\alpha_{p_{n}}$ and $v_{n}$ and $\hat{v}_{n}$ are different minimizers of $\alpha_{n}$ (see Section 3.1). By Theorem 3.4,
$v_{n},\hat{v}_{n}\to c\phi_{1}\quad\text{ in \ }D^{1,2}(\mathbb{R}^{N}),$ (3.25)
where $\phi_{1}$ is the positive first eigenfunction with $|\phi_{1}|_{\infty}=1$ and $c:=(\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{2})^{-1/2}$. By Lemma 3.6 we also know that
$\frac{v_{n}}{|v_{n}|_{\infty}}=\frac{u_{n}}{|u_{n}|_{\infty}}\to\phi_{1}\quad\text{ in \ }C^{1,\alpha}_{loc}(\mathbb{R}^{N}).$ (3.26)
Let $\rho>0$ be such that $\int_{B_{\rho}(0)}|\nabla\phi_{1}|^{2}>0$. Then, by (3.25) and (3.26),
$0=\lim_{n\to\infty}\int_{B_{\rho}(0)}\left|\nabla\left(\frac{v_{n}}{|v_{n}|_{\infty}}-\frac{v_{n}}{c}\right)\right|^{2}=\lim_{n\to\infty}\int_{B_{\rho}(0)}\left|\nabla v_{n}\right|^{2}\left(\frac{1}{|v_{n}|_{\infty}}-\frac{1}{c}\right)^{2}=c^{2}\int_{B_{\rho}(0)}\left|\nabla\phi_{1}\right|^{2}\lim_{n\to\infty}\left(\frac{1}{|v_{n}|_{\infty}}-\frac{1}{c}\right)^{2}.$
We conclude that
[…]
Since $v_{n}\neq\hat{v}_{n}$ by assumption, the function $w_{n}:=\frac{v_{n}-\hat{v}_{n}}{|v_{n}-\hat{v}_{n}|_{\infty}}$ is a nontrivial solution of
$-\Delta w_{n}=Q_{\Omega}\alpha_{n}c_{n}w_{n}=:g_{n}\quad\text{ in }\mathbb{R}^{N}$ (3.28)
where, by the mean value theorem,
$c_{n}(x):=\int_{0}^{1}(p_{n}-1)(tv_{n}(x)+(1-t)\hat{v}_{n}(x))^{p_{n}-2}\,\mathrm{d}t.$
Since $tv_{n}(x)+(1-t)\hat{v}_{n}(x)$ is between $v_{n}(x)$ and $\hat{v}_{n}(x)$, it follows that $c_{n}(x)$ is between $(p_{n}-1)v_{n}(x)^{p_{n}-2}$ and $(p_{n}-1)\hat{v}_{n}(x)^{p_{n}-2}$ for all $x\in\mathbb{R}^{N}$. By (3.27), passing to a subsequence, $|c_{n}|_{\infty}\leq 2$ for all $n\in\mathbb{N}$, $(p_{n}-1)v_{n}^{p_{n}-2}\to 1$, $(p_{n}-1)\hat{v}_{n}^{p_{n}-2}\to 1$ pointwisely in $\mathbb{R}^{N}$, and
$c_{n}(x)\rightarrow 1\qquad\text{ pointwisely in \ }\mathbb{R}^{N}.$
We also know, by Theorem 3.4, that
$\alpha_{n}\to\Lambda_{1}\qquad\text{ as }n\to\infty.$ (3.29)
Then $|g_{n}|_{\infty}\leq C$ and, by elliptic regularity, $w_{n}\to w$ locally uniformly in $\mathbb{R}^{N}$, where $|w|_{\infty}=1$ and $w\neq 0$. Furthermore, testing (3.28) with $w_{n}$,
$\|w_{n}\|^{2}\leq\int_{\mathbb{R}^{N}}Q_{\Omega}\alpha_{n}c_{n}w_{n}^{2}\leq\int_{\Omega}\alpha_{n}c_{n}w_{n}^{2}\leq C|\Omega|.$ (3.30)
Hence, up to a subsequence, $w_{n}$ converges to $w$ weakly in $D^{1,2}(\mathbb{R}^{N})$ and in $L_{loc}^{2}(\mathbb{R}^{N})$. Letting $n\to\infty$ in the weak formulation of (3.28), we obtain that $w$ is a weak solution to
$-\Delta w=Q_{\Omega}\Lambda_{1}(\Omega)w\quad\text{ in }\mathbb{R}^{N},\qquad w>0\quad\text{ in }\mathbb{R}^{N}.$
Then $w=\phi_{1}$. Now, observe that
$0$ $=\int_{\mathbb{R}^{N}}(-\Delta)v_{n}\hat{v}_{n}-(-\Delta)\hat{v}_{n}v_{n}=\int_{\mathbb{R}^{N}}Q_{\Omega}\alpha_{n}v_{n}\hat{v}_{n}(v_{n}^{p_{n}-2}-\hat{v}_{n}^{p_{n}-2})$
$=\int_{\mathbb{R}^{N}}Q_{\Omega}\alpha_{n}v_{n}\hat{v}_{n}\left(\int_{0}^{1}(sv_{n}+(1-s)\hat{v}_{n})^{p_{n}-3}\,\mathrm{d}s\right)(p_{n}-2)(v_{n}-\hat{v}_{n}).$
This implies that
$0$ $=\int_{\mathbb{R}^{N}}Q_{\Omega}v_{n}\hat{v}_{n}w_{n}d_{n},\qquad\text{where \ }d_{n}:=\int_{0}^{1}(sv_{n}+(1-s)\hat{v}_{n})^{p_{n}-3}\,\mathrm{d}s.$
By (3.25), $sv_{n}+(1-s)\hat{v}_{n}\to c\phi_{1}$ pointwisely in $\mathbb{R}^{N}.$ Moreover, by Lemma 3.7, (3.26), and (3.27),
$|Q_{\Omega}v_{n}\hat{v}_{n}w_{n}d_{n}|$ $\leq v_{n}\hat{v}_{n}\min\{v_{n}^{p_{n}-3},\hat{v}_{n}^{p_{n}-3}\}=\min\{v_{n}^{p_{n}-2}\hat{v}_{n},v_{n}\hat{v}_{n}^{p_{n}-2}\}$
$\leq\max\{|v_{n}|_{\infty}^{p_{n}-2},|\hat{v}_{n}|_{\infty}^{p_{n}-2}\}\min\left\{|\hat{v}_{n}|_{\infty}\frac{\hat{v}_{n}}{|\hat{v}_{n}|_{\infty}},|v_{n}|_{\infty}\frac{v_{n}}{|v_{n}|_{\infty}}\right\}$
$\leq C\min\left\{\frac{\hat{u}_{n}}{|u_{n}|_{\infty}},\frac{u_{n}}{|u_{n}|_{\infty}}\right\}\leq C\min\{1,|x|^{-\frac{2}{q-2}}\}$
for some $C>0$, $n$ large enough, and $q>2$ so that $x\mapsto|x|^{-\frac{2}{q-2}}\in L^{1}(\mathbb{R}^{N}\backslash B_{1}(0))$. Then, by dominated convergence,
$0$ $=\lim_{n\to\infty}\int_{\mathbb{R}^{N}}Q_{\Omega}v_{n}\hat{v}_{n}w_{n}d_{n}=c^{\prime}\int_{\mathbb{R}^{N}}Q_{\Omega}\phi_{1}^{2}>0$
for some $c^{\prime}>0,$ which is a contradiction.
Remark 3.8.
In the classical blow-up argument to show uniqueness of positive solutions, as in [20], the idea is to consider two different positive solutions $u_{n}$ and $\hat{u}_{n}$, to define $z_{n}:=\frac{u_{n}-\hat{u}_{n}}{|u_{n}-\hat{u}_{n}|_{\infty}}$ and to show that $z_{n}\to\phi_{1}$ locally. This can be done also in the setting of problem (1.1). However, the next steps of the arguments are problematic. One wishes to show that $z_{n}$ must change sign. This is usually done by contradiction and using integration by parts, but in our case we obtain that
$0=\int_{\mathbb{R}^{N}}(-\Delta)u_{n}\hat{u}_{n}-u_{n}(-\Delta)\hat{u}_{n}=\int_{\mathbb{R}^{N}}Q_{\Omega}u_{n}\hat{u}_{n}(u_{n}^{p_{n}-2}-\hat{u}_{n}^{p_{n}-2}).$
Due to the weight $Q_{\Omega}$, this does not yield a direct contradiction and one cannot conclude as in the classical case with Dirichlet boundary conditions (and without the coefficient $Q_{\Omega}$).
4 Sharp decay below and at the Serrin exponent
Let $N\geq 3$ and let $p_{s}:=\frac{2N-2}{N-2}$ denote the Serrin exponent. Next we show Proposition 1.5.
Proof of Proposition 1.5.
[…]
$-\Delta z=-w^{p-1}+Cv^{p-1}+((Cv)^{p-1}-(Cv)^{p-1})=-w^{p-1}+(Cv)^{p-1}+(1-C^{p-2})Cv^{p-1}\quad\text{for $x\in\mathbb{R}^{N}\backslash\Omega$}.$
Let
$c(x):=\begin{cases}\frac{w^{p-1}(x)-(Cv(x))^{p-1}}{w(x)-Cv(x)},&\text{if \ }w(x)\neq Cv(x),\\ 0,&\text{if \ }w(x)=Cv(x).\end{cases}$ (4.1)
Then $c\geq 0$, and $z$ is a classical solution to
$-\Delta z+cz=(1-C^{p-2})Cv^{p-1}\geq 0\quad\text{ in }\mathbb{R}^{N}\backslash\Omega,\qquad z>0\quad\text{ on }\partial\Omega.$
By the maximum principle, $z\geq 0$ in $\mathbb{R}^{N}\backslash\Omega$. Indeed, assume by contradiction that $x_{0}\in\mathbb{R}^{N}\backslash\Omega$ is such $z(x_{0})<0$ (note that $\lim_{|x|\to\infty}z(x)=0$). Then $\Delta z(x_{0})\leq 0$, $c(x_{0})z(x_{0})<0$ and, therefore,
$0>-\Delta z(x_{0})+c(x_{0})z(x_{0})\geq 0,$
which would yield a contradiction. Then, since $\Omega$ is a compact set, there exists $C^{\prime}>0$ such that $w(x)\geq C^{\prime}|x|^{-\frac{2}{p-2}}$ for all $|x|>1$.
Now we obtain the bounds from above. Using the same notation, if $C>1$ is such that $z<0$ on $\partial\Omega.$ Then,
$-\Delta z+cz=(1-C^{p-2})Cv^{p-1}\leq 0\quad\text{ in }\mathbb{R}^{N}\backslash\Omega,\qquad z<0\quad\text{ on }\partial\Omega.$
Arguing as before, $z\leq 0$ in $\mathbb{R}^{N}\backslash\Omega$, namely, there exists $C^{\prime}>0$ such that $w(x)\leq C^{\prime}|x|^{-\frac{2}{p-2}}$ for $|x|>1.$ This yields (1.6).
Case $p=p_{s}$: This case can be argued similarly using a different comparison function. Let
$v(x)=k_{N}\left(|x|\sqrt{\ln(|x|)}\right)^{2-N},\qquad k_{N}:=2^{2-N}\left(3N^{2}-10N+8\right)^{\frac{N-2}{2}}.$
Then, by direct calculations,
$-\Delta v+v^{\frac{N}{N-2}}=c_{N}\frac{\ln(r)-1}{r^{N}(\ln(r))^{\frac{N+2}{2}}}>0\quad\text{ in }\mathbb{R}^{N}\backslash\overline{B_{1}(0)}$
with $c_{N}:=2^{-N}(N-2)^{N/2}N(3N-4)^{\frac{N-2}{2}}.$ Let z:=wCvz:=w-Cv, with $C>1$ such that $z<0$ on $\partial B_{R}(0)$ with $R>1$ such that $\Omega\subset B_{R}(0).$ Then,
$-\Delta z<-w^{\frac{N}{N-2}}+Cv^{\frac{N}{N-2}}+((Cv)^{\frac{N}{N-2}}-(Cv)^{\frac{N}{N-2}})=-w^{\frac{N}{N-2}}+(Cv)^{\frac{N}{N-2}}+(1-C^{\frac{2}{N-2}})Cv^{\frac{N}{N-2}}\quad\text{for $x\in\mathbb{R}^{N}\backslash B_{R}(0)$}.$
Let $c$ as in (4.1). Then $c\geq 0$ and $z$ is a classical solution to
$-\Delta z+cz<(1-C^{\frac{2}{N-2}})Cv^{\frac{N}{N-2}}<0\quad\text{ in }\mathbb{R}^{N}\backslash B_{R}(0),\qquad z<0\quad\text{ on }\partial B_{R}(0).$
As before, by the maximum principle, $z\leq 0$ in $\mathbb{R}^{N}\backslash B_{R}(0)$. Since $\Omega$ is a compact set, there exists $C^{\prime}>0$ such that $w(x)\leq C^{\prime}\left(|x|\sqrt{\ln(|x|)}\right)^{2-N}$ for all $|x|>1$.
Finally, we argue the bound from below in (1.7). For this, we must adjust the constants. Let
$V(x)=K_{N}\left(|x|\sqrt{\ln(|x|)}\right)^{2-N},\qquad K_{N}:=2^{\frac{2-N}{2}}(N-2)^{N-2}.$
Then, by direct calculations,
$-\Delta V+V^{\frac{N}{N-2}}=C_{N}\frac{-1}{r^{N}(\ln(r))^{\frac{N+2}{2}}}<0\quad\text{ in }\mathbb{R}^{N}\backslash\overline{B_{1}(0)}$
with $C_{N}:=2^{-\frac{N+2}{2}}(N-2)^{N-1}N>0$. Let $Z:=w-\varepsilon V$, with $\varepsilon\in(0,1)$ such that $Z>0$ on $\partial B_{R}(0)$ with $R>1$ such that $\Omega\subset B_{R}(0).$ Then,
$-\Delta Z>-w^{\frac{N}{N-2}}+\varepsilon V^{\frac{N}{N-2}}+((\varepsilon V)^{\frac{N}{N-2}}-(\varepsilon V)^{\frac{N}{N-2}})=-w^{\frac{N}{N-2}}+(\varepsilon V)^{\frac{N}{N-2}}+(1-\varepsilon^{\frac{2}{N-2}})\varepsilon V^{\frac{N}{N-2}}\quad\text{for $x\in\mathbb{R}^{N}\backslash B_{R}(0)$}.$
Hence,
$-\Delta Z+cZ\geq(1-\varepsilon^{\frac{2}{N-2}})\varepsilon V^{\frac{N}{N-2}}>0\quad\text{ in }\mathbb{R}^{N}\backslash B_{R}(0),\qquad Z>0\quad\text{ on }\partial B_{R}(0).$
Arguing as before, the maximum principle yields that $Z\geq 0$ in $\mathbb{R}^{N}\backslash B_{R}(0)$ and, since $w>0$ in $\mathbb{R}^{N},$ we can find $\varepsilon^{\prime}>0$ such that $w(x)>\varepsilon^{\prime}\left(|x|\sqrt{\ln(|x|)}\right)^{2-N}$ for $|x|>1.$
5 Multiplicity of least-energy solutions
In this section we show a symmetry breaking and multiplicity result for least energy solutions. Let us consider the equation
$\begin{cases}-\Delta u=Q_{\Omega}|u|^{2^{\ast}-2}u,\\ u\in D^{1,2}(\mathbb{R}^{N}).\end{cases}$ (5.1)
This problem is studied in [5]. The solutions of (5.1) are the critical points of the functional $J_{\ast}:D^{1,2}(\mathbb{R}^{N})\to\mathbb{R}$ given by
$J_{\ast}(u):=\frac{1}{2}\|u\|^{2}-\frac{1}{2^{\ast}}\int_{\mathbb{R}^{N}}Q_{\Omega}(x)|u|^{2^{\ast}},\qquad v\in D^{1,2}(\mathbb{R}^{N}).$
The nontrivial critical points of $J_{\ast}$ belong to the Nehari manifold
$\mathcal{N}_{\ast}=\Big\{u\in D^{1,2}(\mathbb{R}^{N})\setminus\{0\}\,:\,\|u\|^{2}=\int_{\mathbb{R}^{N}}Q_{\Omega}(x)|u|^{2^{\ast}}\Big\}.$
Lemma 5.1.
For $p\in(2,2^{\ast})$, let $u_{p}$ be a positive least energy solution of (1.1). There exists $p_{0}\in(2,2^{\ast})$ such that $\int_{\mathbb{R}^{N}}Q_{\Omega}(x)|u_{p}|^{2^{\ast}}>0$ for all $p\in(p_{0},2^{\ast})$. Let $t_{p}\in(0,\infty)$ be such that $\tilde{u}_{p}:=t_{p}u_{p}\in\mathcal{N}_{\ast}$, i.e.,
$t_{p}=\left(\frac{\|u_{p}\|^{2}}{\int_{\mathbb{R}^{N}}Q_{\Omega}(x)|u_{p}|^{2^{\ast}}}\right)^{\frac{1}{2^{\ast}-2}}=\left(\frac{\int_{\mathbb{R}^{N}}Q_{\Omega}(x)|u_{p}|^{p}}{\int_{\mathbb{R}^{N}}Q_{\Omega}(x)|u_{p}|^{2^{\ast}}}\right)^{\frac{1}{2^{\ast}-2}}.$ (5.2)
Then,
$\lim_{p\to 2^{\ast}}t_{p}=1\qquad\text{and}\qquad\lim_{p\to 2^{\ast}}J_{\ast}(\tilde{u}_{p})=c_{\ast}:=\inf_{u\in\mathcal{N}_{\ast}}J_{\ast}(u)=\frac{1}{N}S^{\frac{N}{2}},$
where $S$ is the best constant for the Sobolev embedding $D^{1,2}(\mathbb{R}^{N})\hookrightarrow L^{2^{*}}(\mathbb{R}^{N}).$
Proof.
By [5, Proof of Proposition 2.3] we have that
$\alpha_{2^{*}}=\inf_{\begin{subarray}{c}u\in D^{1,2}(\mathbb{R}^{N})\\ \int_{\mathbb{R}^{N}}Q_{\Omega}|u|^{2^{*}}>0\end{subarray}}\frac{\|u\|^{2}}{\left(\displaystyle\int_{\mathbb{R}^{N}}Q_{\Omega}|u|^{2^{*}}\right)^{2/2^{*}}}=\inf_{\begin{subarray}{c}u\in D_{0}^{1,2}(\Omega)\\ u\neq 0\end{subarray}}\frac{\displaystyle\int_{\Omega}|\nabla u|^{2}}{\left(\displaystyle\int_{\Omega}|u|^{2^{*}}\right)^{2/2^{*}}}=\inf_{\begin{subarray}{c}u\in D^{1,2}(\mathbb{R}^{N})\\ u\neq 0\end{subarray}}\frac{\|u\|^{2}}{\left(\displaystyle\int_{\mathbb{R}^{N}}|u|^{2^{*}}\right)^{2/2^{*}}}.$ (5.3)
Moreover, for $u\in D^{1,2}(\mathbb{R}^{N})$,
$\left(\int_{\mathbb{R}^{N}}Q_{\Omega}|u|^{p}\right)^{\frac{1}{p}}\leq\left(\int_{\Omega}|u|^{p}\right)^{\frac{1}{p}}\leq|\Omega|^{\frac{2^{*}-p}{2^{*}p}}\left(\int_{\Omega}|u|^{2^{*}}\right)^{\frac{1}{2^{*}}}\leq|\Omega|^{\frac{2^{*}-p}{2^{*}p}}\left(\int_{\mathbb{R}^{N}}|u|^{2^{*}}\right)^{\frac{1}{2^{*}}}.$ (5.4)
Hence,
$\alpha_{2^{*}}$ $=\inf_{\begin{subarray}{c}u\in D^{1,2}(\mathbb{R}^{N})\\ u\neq 0\end{subarray}}\frac{\|u\|^{2}}{\left(\displaystyle\int_{\mathbb{R}^{N}}|u|^{2^{*}}\right)^{2/2^{*}}}\leq\inf_{\begin{subarray}{c}u\in D^{1,2}(\mathbb{R}^{N})\\ \int_{\mathbb{R}^{N}}Q_{\Omega}|u|^{p}>0\end{subarray}}\frac{\|u\|^{2}}{\left(\displaystyle\int_{\mathbb{R}^{N}}|u|^{2^{*}}\right)^{2/2^{*}}}$
$\leq|\Omega|^{\frac{2^{*}-p}{2^{*}p}}\inf_{\begin{subarray}{c}u\in D^{1,2}(\mathbb{R}^{N})\\ \int_{\mathbb{R}^{N}}Q_{\Omega}|u|^{p}>0\end{subarray}}\frac{\|u\|^{2}}{\left(\displaystyle\int_{\mathbb{R}^{N}}Q_{\Omega}|u|^{p}\right)^{2/p}}=|\Omega|^{\frac{2^{*}-p}{2^{*}p}}\alpha_{p}.$ (5.5)
By Lemma 3.1, passing to the limit as $p\to 2^{\ast}$, we have that
$\alpha_{2^{\ast}}\leq\liminf_{p\to 2^{\ast}}\alpha_{p}\leq\limsup_{p\to 2^{\ast}}\alpha_{p}\leq a_{1}<\infty.$
Now, we claim that $\lim\limits_{p\to 2^{*}}\alpha_{p}=\alpha_{2^{*}}.$ Arguing by contradiction, assume that
$\alpha_{2^{\ast}}+\varepsilon<\limsup\limits_{p\to 2^{\ast}}\alpha_{p}\qquad\text{ for some $\varepsilon>0$.}$ (5.6)
From (5.3) and the density of $C_{c}^{\infty}(\mathbb{R}^{N})$ in $D^{1,2}(\mathbb{R}^{N})$, there exists $\tilde{u}\in C_{c}^{\infty}(\mathbb{R}^{N})$ such that
$\frac{\|\tilde{u}\|^{2}}{\left(\int_{\mathbb{R}^{N}}Q_{\Omega}|\tilde{u}|^{2^{\ast}}\right)^{2/2^{\ast}}}<\alpha_{2^{\ast}}+\frac{\varepsilon}{2}.$ (5.7)
Since the function $p\mapsto\int_{\mathbb{R}^{N}}Q_{\Omega}|\tilde{u}|^{p}$ is continuous in $(2,2^{*}]$, there exists $q\in(2,2^{\ast})$ such that
$0<\int_{\mathbb{R}^{N}}Q_{\Omega}|\tilde{u}|^{p}\qquad\text{and}\qquad\left|\frac{\|\tilde{u}\|^{2}}{\left(\int_{\mathbb{R}^{N}}Q_{\Omega}|\tilde{u}|^{2^{\ast}}\right)^{2/2^{\ast}}}-\frac{\|\tilde{u}\|^{2}}{\left(\int_{\mathbb{R}^{N}}Q_{\Omega}|\tilde{u}|^{p}\right)^{2/p}}\right|<\frac{\varepsilon}{2}\qquad\text{ for all }\,p\in[q,2^{\ast}).$ (5.8)
Then, by (5.7), (5.8), and (5.6),
$\alpha_{p}\leq\frac{\|\tilde{u}\|^{2}}{\left(\int_{\mathbb{R}^{N}}Q_{\Omega}|\tilde{u}|^{p}\right)^{2/p}}<\alpha_{2^{\ast}}+\varepsilon<\limsup_{p\to 2^{\ast}}\alpha_{p}\qquad\text{ for all }\,p\in[q,2^{\ast}),$
which yields a contradiction. Hence,
$\alpha_{2^{\ast}}=\lim_{p\to 2^{\ast}}\alpha_{p}.$ (5.9)
Let $u_{p}$ be a positive minimizer of (1.1). It is easy to see that
$\frac{\|u_{p}\|^{2}}{\left(\int_{\mathbb{R}^{N}}Q_{\Omega}(x)|u_{p}|^{p}\right)^{2/p}}=\alpha_{p}.$ (5.10)
Then, by (5.4),
$\alpha_{2^{\ast}}\leq\frac{\|u_{p}\|^{2}}{\left(\int_{\mathbb{R}^{N}}|u_{p}|^{2^{\ast}}\right)^{2/2^{\ast}}}\leq|\Omega|^{\frac{2(2^{\ast}-p)}{2*p}}\frac{\|u_{p}\|^{2}}{\left(\int_{\mathbb{R}^{N}}Q_{\Omega}(x)|u_{p}|^{p}\right)^{2/p}}=|\Omega|^{\frac{2(2^{\ast}-p)}{2*p}}\alpha_{p}.$
Passing to the limit as $p\to 2^{*}$ and using (5.9),
$\lim_{p\to 2^{\ast}}\int_{\mathbb{R}^{N}}|u_{p}|^{2^{\ast}}=\lim_{p\to 2^{\ast}}\int_{\mathbb{R}^{N}}Q_{\Omega}|u_{p}|^{p}.$ (5.11)
From (5.4), (5.11), and (5.10) (using that $\|u_{p}\|^{2}=\int_{\mathbb{R}^{N}}Q_{\Omega}|u_{p}|^{p}$),
$\lim_{p\to 2^{\ast}}\int_{\Omega}|u_{p}|^{p}=\lim_{p\to 2^{\ast}}\int_{\Omega}|u_{p}|^{2^{\ast}}=\lim_{p\to 2^{\ast}}\int_{\mathbb{R}^{N}}|u_{p}|^{2^{\ast}}=\lim_{p\to 2^{\ast}}\int_{\mathbb{R}^{N}}Q_{\Omega}|u_{p}|^{p}=\lim_{p\to 2^{\ast}}\alpha_{p}^{\frac{p}{p-2}}=\alpha_{2^{\ast}}^{\frac{N}{2}}.$ (5.12)
This implies that $\lim_{p\to 2^{\ast}}\int_{\mathbb{R}^{N}\setminus\Omega}|u_{p}|^{2^{\ast}}=0$ and
$\lim_{p\to 2^{\ast}}\int_{\mathbb{R}^{N}}Q_{\Omega}|u_{p}|^{2^{\ast}}=\alpha_{2^{\ast}}^{\frac{N}{2}}.$ (5.13)
Therefore, $\int_{\mathbb{R}^{N}}Q_{\Omega}|u_{p}|^{2^{\ast}}>0$ for $p$ close enough to $2^{\ast}$. The natural projection of $u_{p}$ to the Nehari manifold $\mathcal{N}_{2^{\ast}}$ is given by $\widehat{u}_{p}=t_{p}u_{p}$ where $t_{p}$ is defined in (5.2). By (5.12) and (5.13), $\lim\limits_{p\to 2^{\ast}}t_{p}=1$.
Finally, note that
$J_{\ast}(\widehat{u}_{p})=\frac{2^{\ast}-2}{2^{\ast}2}\|\widehat{u}_{p}\|^{2}=\frac{1}{N}t_{p}^{2}\|u_{p}\|^{2}=\frac{1}{N}t_{p}^{2}\alpha_{p}^{\frac{p}{p-2}}.$
Hence, $\lim_{p\to 2^{\ast}}J_{\ast}(\widehat{u}_{p})=\frac{1}{N}\alpha_{2^{\ast}}^{\frac{N}{2}}=c_{\ast},$ as claimed. That $c_{\ast}=\frac{1}{N}S^{\frac{N}{2}}$ follows from [5, Proposition 2.3 (c)]. ∎
Corollary 5.2.
Let $p_{k}\in(2,2^{\ast})$ be such that $p_{k}\to 2^{\ast}$ and let $u_{p_{k}}$ be a positive minimizer to (1.1) with $p=p_{k}$. Then, after passing to a subsequence, there exist $\varepsilon_{k}>0$ and $\xi_{k}\in\Omega$ such that
$\lim_{k\to\infty}\varepsilon_{k}^{-1}\text{dist}(\xi_{k},\partial\Omega)=\infty\qquad\text{and}\qquad\lim_{k\to\infty}\|u_{k}-U_{\varepsilon_{k},\xi_{k}}\|=0,$
where
$U_{\varepsilon_{k},\xi_{k}}(x):=(N(N-2))^{\frac{N-2}{4}}\frac{\varepsilon_{k}^{\frac{N-2}{2}}}{\left(\varepsilon_{k}^{2}+|x-\xi_{k}|^{2}\right)^{\frac{N-2}{2}}}.$
Proof.
Let $u_{p_{k}}$ be as in the statement and let $\widehat{u}_{p_{k}}=t_{p_{k}}u_{p_{k}}\in\mathcal{N}_{\ast}$ with $t_{p_{k}}$ as in Lemma 5.1. By [5, Corollary 4.4] and Lemma 5.1, there exist $\varepsilon>0$ and $\xi_{k}\in\Omega$ such that $\lim\limits_{k\to\infty}\varepsilon_{k}^{-1}\operatorname{dist}(\xi_{k},\partial\Omega)=\infty$ and $\lim\limits_{k\to\infty}\|\widehat{u}_{p_{k}}-U_{\varepsilon_{k},\xi_{k}}\|=0.$ Hence,
$\|u_{p_{k}}-U_{\varepsilon_{k},\xi_{k}}\|\leq\frac{1}{t_{p_{k}}}\Big(\|\tilde{u}_{p_{k}}-U_{\varepsilon_{k},\xi_{k}}\|+|1-t_{p_{k}}|\|U_{\varepsilon_{k},\xi_{k}}\|\Big)=o(1)\qquad\text{ as }k\to\infty,$
and the statement follows. ∎
We now use Corollary 5.2 to show the symmetry breaking result for least energy solutions of (1.1) for $p$ close to $2^{\ast}$ in annuli stated in the Introduction. As a consequence, one can use rotations of this solution to obtain infinitely many different least energy solutions of (1.1). This contrasts with the case of $p$ close to 2, where, by Theorem 1.1, the least energy solution is unique (and therefore radially symmetric whenever $\Omega$ is invariant under rotations).
Proof of Theorem 1.2.
Let $p_{k}\in(2,2^{\ast})$ be such that $p_{k}\to 2^{\ast}$ and let $u_{p_{k}}$ be a positive least energy solution of (1.1) with $p=p_{k}$. By Corollary 5.2, there are sequences $(\varepsilon_{k})\subset(0,\infty)$ and $(\xi_{k})\subset\Omega$ such that, passing to a subsequence,
$u_{p_{k}}=U_{\varepsilon_{k},\xi_{k}}+o(1)\quad\text{ in $D^{1,2}(\mathbb{R}^{N})$},\qquad\varepsilon_{k}\to 0,\qquad\xi_{k}\to\xi_{0}\in\overline{\Omega},$
as $k\to\infty$. The result now follows since $U_{\varepsilon_{k},\xi_{k}}$ is not invariant under rotations with respect to the origin. The foliated Schwarz symmetry follows from [6, Theorem 1.4]. ∎
Acknowledgments
We thank Bendetta Pellaci for helpful discussions.